### v2

# Dissipative solutions to Stochastic 3D Euler equations

Abstract.
We construct probabilistically strong solutions to the three-dimensional Euler equations perturbed by additive noise that are $\mathbb{P}$-almost surely continuous in time, Hölder in space, and satisfy the local energy inequality up to an arbitrarily large stopping time.
Key words and phrases:
Stochastic Euler equations, Local energy inequality, Convex integration
1. Introduction
In this work, we consider the stochastic Euler equations on the three-dimensional torus $\mathbb{T}^{3}:=\mathbb{R}^{3}/2\pi\mathbb{Z}^{3}$ driven by additive noise
(SE) $\begin{cases}du+\mathord{{\rm div}}(u\otimes u)\,dt+\nabla p\,dt=Q^{1/2}dW,\\ \mathord{{\rm div}}(u)=0.\end{cases}$
Euler equations govern the evolution of an ideal fluid and constitute one of the most important models in fluid dynamics. The presence of a random external force in (SE) can help represent many phenomena perturbing the dynamics that are neglected by the deterministic equation, such as endogenous microscopic thermal effects or exogenous influences of the environment on the fluid. First investigations on (SE) date back to the works [MiVa00, Ki09], where local existence of smooth solutions is proved (see also [BeFl99] for other results in dimension two). Measure-valued solutions satisfying weak-strong uniqueness and a version of the energy inequality have been constructed in [BrMo21], building upon results by DiPerna and Majda [DPMa87] for the deterministic Euler equations $Q=0$. Whether solutions are unique or not remained an open problem until the work [HoZhZh22] by Hofmanová, Zhu, and Zhu, adapting for the first time the convex integration scheme developed by De Lellis and Székelyhidi Jr. in [DLSz09] to the stochastic case, and obtaining non-uniqueness in law for (SE). After [HoZhZh22] made clear that convex integration techniques can be applied to stochastic equations (let us also mention [ChFeFl21] in this regard), a sequence of papers has appeared [HoZhZh25, LuZh24, KiKo24, LuLuZh25] extending the works [DLS13, DLS14, BuDLIsSz15, BuDLIsSz16, Is18, BuDLSzVi19] on the deterministic Euler equations. In particular, for every $\vartheta\in(0,1/3)$, Lü, Lü, and Zhu proved in [LuLuZh25] existence and non-uniqueness of solutions of (SE) with regularity $u\in C([0,\infty),C^{\vartheta}(\mathbb{T}^{3},\mathbb{R}^{3})$ almost surely, which are strictly dissipative up to an arbitrary large stopping time $\mathfrak{t}$ in the sense that $u$ satisfies almost surely the energy inequality
(1.1) $\frac{1}{2}\|u(t\wedge\mathfrak{t})\|_{L^{2}_{x}}^{2}\leq\frac{1}{2}\|u(s\wedge\mathfrak{t})\|_{L^{2}_{x}}^{2}+\int_{s\wedge\mathfrak{t}}^{t\wedge\mathfrak{t}}\langle u(r),Q^{1/2}dW_{r}\rangle+\frac{1}{2}\mbox{Tr}(Q)(t\wedge\mathfrak{t}-s\wedge\mathfrak{t}),$
for every $s<t$, and the inequality is strict if $s<\mathfrak{t}$. In the present work we investigate solutions of (SE) that satisfy almost surely the local energy inequality up to time $\mathfrak{t}$:
(LEI) $\frac{1}{2}d|u|^{2}+\mathord{{\rm div}}\left(\left(\frac{1}{2}|u|^{2}+p\right)u\right)dt-\mathfrak{q}\,dt-u\cdot Q^{1/2}dW\leq 0,$
where $\mathfrak{q}:=\frac{1}{2}\sum_{k}|Q^{1/2}e_{k}|^{2}$ for a CONS $\{e_{k}\}_{k\in\mathbb{N}}$ of the space of $L^{2}_{x}$ solenoidal vector fields, $d|u|^{2}$ and $dW$ are Itō stochastic differentials, and the equation is meant as an almost sure identity in the sense of distributions. Notice that (1.1) can be obtained from (LEI) by space integration.
The importance of a local form of energy inequality towards partial regularity results for suitable weak solutions of the Navier-Stokes equations is known since the works of Scheffer [Sc77] and Caffarelli, Kohn, and Niremberg [CaKoNi82]. Similar results are available for the stochastic Navier-Stokes equations, see [Ro10, FlRo02]. The local energy balance (LEI) is a natural condition to impose also on the stochastic Euler equations (SE), since any physically relevant $L^{3}_{t,x}$ zero-viscosity limit of suitable solutions of the stochastic Navier-Stokes equations must inherit this property. It describes the local change in kinetic energy due to spatial roughness of $u$: indeed, the equality in (LEI) holds for smooth solutions $u$, and in general the inequality corresponds to no local creation of kinetic energy due to irregularities of $u$. We refer to the celebrated work of Duchon and Robert [DuRo00] for more details.
In the deterministic setting $Q=0$, existence and non-uniqueness of solutions to the system (SE)-(LEI) have been shown in the series of papers [DLSz10, Is22, DLK23, GiKwNo]. The purpose of the present work is to provide similar results in the stochastic case $Q\neq 0$. More specifically, we will prove the following:
Theorem 1.1.
Let $(\Omega,\mathbb{F},\{\mathbb{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a stochastic basis with right-continuous and complete filtration supporting a sufficiently smooth-in-space $Q$-Wiener process $Q^{1/2}W$ with zero space average and null divergence. Let be given $T>0$, $\varkappa\in(0,1)$, $\alpha\in(0,1/7)$ and $E\in C^{1}_{loc}([0,\infty),\mathbb{R})$. Then there exist two $\{\mathbb{F}_{t}\}_{t\geq 0}$-progressively measurable stochastic processes
(1.2) $u:\Omega\to C([0,\infty),C^{\alpha}(\mathbb{T}^{3},\mathbb{R}^{3})),\qquad p:\Omega\to C([0,\infty),C^{2\alpha}(\mathbb{T}^{3},\mathbb{R})),$
such that $(u,p)$ is $\mathbb{P}$-almost surely a solution of (SE) on the time interval $0,\infty)$ in analytically weak sense, and a stopping time $\mathfrak{t}$ satisfying $\mathbb{P}\{\mathfrak{t}\geq T\}\geq\varkappa$ such that the local energy equality
(LEE) $\frac{1}{2}d|u|^{2}+\mathord{{\rm div}}\left(\left(\frac{1}{2}|u|^{2}+p\right)u\right)dt-\mathfrak{q}\,dt-u\cdot Q^{1/2}dW=-E^{\prime}dt$
holds $\mathbb{P}$-almost surely in the sense of distributions on the time interval $0,\mathfrak{t}]$. In particular, (LEI) holds strictly if $E^{\prime}>0$. Finally, non-uniqueness in law for $(u,p)$ holds within this class.
Acknowledgements
UP has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 949981).
FT is supported by the Istituto Nazionale di Alta Matematica (INdAM) through the project GNAMPA 2025 “Modelli stocastici in Fluidodinamica e Turbolenza” -CUP E5324001950001. FT acknowledges the hospitality of Universität Bielefeld, where part of this work has been done.
2. Preliminaries and main iterative proposition
2.1. Reformulation of the problem without local martingales
It is convenient to rewrite (SE)-(LEE) without the local martingale terms $Q^{1/2}dW$ and $u\cdot Q^{1/2}dW$. To attain this, we reformulate the problem using the so-called Da Prato-Debussche trick, namely we decompose solutions of (SE) as $u=v+z$, where $z:=Q^{1/2}W$ and $v$ solves the random PDE
(2.1) $\begin{cases}\partial_{t}v+\mathord{{\rm div}}\big((v+z)\otimes(v+z)\big)+\nabla p=0,\\ \mathord{{\rm div}}(v)=0.\end{cases}$
A local energy inequality can be derived for $v$ as follows. First, rewrite the stochastic differential
(2.2) $\frac{1}{2}d|v|^{2}=\frac{1}{2}d|u-z|^{2}=\frac{1}{2}d|u|^{2}+\frac{1}{2}d|z|^{2}-d(u\cdot z).$
Since by assumption $z$ is spatially smooth enough, Itō Formula gives
(2.3) $\frac{1}{2}d|z|^{2}=z\cdot dz+\mathfrak{q}\,dt,\quad d(u\cdot z)=du\cdot z+u\cdot dz+2\mathfrak{q}\,dt,$
so that the local energy equality (LEE) is equivalent to
(2.4) $\frac{d}{dt}\frac{|v|^{2}}{2}+v\cdot\nabla p+v\cdot\mathord{{\rm div}}\big((v+z)\otimes(v+z)\big)=-E^{\prime}.$
We refer to [Ro10] for additional details. Therefore, solutions $(u,p)$ of (SE)-(LEE) are equivalent to solutions of the following system:
(2.5) $\begin{cases}\partial_{t}v+\mathord{{\rm div}}\big((v+z)\otimes(v+z)\big)+\nabla p=0,\\ \mathord{{\rm div}}(v)=0,\\ \partial_{t}\frac{|v|^{2}}{2}+v\cdot\nabla p+v\cdot\mathord{{\rm div}}\big((v+z)\otimes(v+z)\big)=-E^{\prime}.\end{cases}$
2.2. Notation
Let $\mathbb{N}$ denote the set of natural numbers, including 0 in our convention, and let $n,m\in\mathbb{N}$. For a multi-index $\mathbf{r}=(r_{1},r_{2},\dots,r_{n})\in\mathbb{N}^{n}$ of order $|\mathbf{r}|:=r_{1}+r_{2}+\dots+r_{n}$ and a smooth function $f:\mathcal{O}\to\mathbb{R}^{m}$ defined on a smooth domain $\mathcal{O}\subset\mathbb{R}^{n}$ without boundary, we denote $D^{\mathbf{r}}f:=\partial_{1}^{r_{1}}\partial_{2}^{r_{2}}\dots\partial_{n}^{r_{n}}f$. With a slight abuse of notation, in the following we identify functions defined on the torus $\mathbb{T}^{3}$ as periodic functions defined on $\mathcal{O}=\mathbb{R}^{3}$.
Let $N\in\mathbb{N}$ and $\delta\in(0,1)$, and let us define the following Hölder seminorms:
(2.6) $f]_{N}:=\max\limits_{|\mathbf{r}|=N}\|D^{\mathbf{r}}f\|_{C^{0}},\quad[f]_{N+\delta}:=\max\limits_{|\mathbf{r}|=N}\sup\limits_{x\neq y}\frac{|D^{\mathbf{r}}f(x)-D^{\mathbf{r}}f(y)|}{|x-y|^{\delta}}.$
Then, the Hölder norms are defined as
(2.7) $\|f\|_{N}:=\max_{j=0,\dots,N}\{[f]_{j}\},\quad\|f\|_{N+\delta}:=\max\{\|f\|_{N},[f]_{N+\delta}\}.$
Furthermore, in order to simplify the statement of many propositions here contained, we adopt the notational convention $\|f\|_{M}\equiv 0$ whenever the index $M<0$.
2.3. Preparation of the convex integration scheme
The proof of ˜1.1 relies on a convex integration scheme, mostly inspired by the works [Is22, DLK23, LuLuZh25]. Let us introduce various parameters that will allow to estimate all the objects appearing in the upcoming sections. Define for $q\in\mathbb{N}$:
(2.8) $\lambda_{q}:=\lfloor a^{b^{q}}\rfloor,\quad\overline{\delta}_{q}:=\lambda_{q}^{-2\alpha},\quad\delta_{q}:=\begin{cases}\lambda_{0}^{\gamma},\,&q=0,\\ \lambda_{0}^{\gamma}\overline{\delta}_{q}\lambda_{1}^{2\alpha},\,&q\geq 1,\end{cases}$
where $a>2$ will be very large, $b>1$ will be chosen close to 1 and $\alpha\in(0,1/7)$ indicates the Hölder regularity of the solution identified by the iterative procedure.
Let also also denote $N_{1}=n_{0}+5$, where the parameter $n_{0}\in\mathbb{N}$ will be chosen in the following sections and will depend on the value $b$. In particular, $n_{0}$ gets larger as $b$ is closer to $1$. We assume that the $Q$-Wiener process $z$ takes values $\mathbb{P}$-almost surely in the Sobolev space $H^{2+N_{1}}(\mathbb{T}^{3},\mathbb{R}^{3})$ with zero mean and divergence.
Let us also introduce $0<\delta\ll 1$ sufficiently small such that
(2.9) $\frac{1/2+\delta}{1/2-\delta}<\frac{1-4\alpha}{3\alpha}\wedge\frac{4}{3},$
and $L>1$ such that the following stopping time
(2.10) $\mathfrak{t}:=\inf\left\{t\geq 0:\|z\|_{C^{1/2-\delta}([0,t],C_{x}^{N_{1}})}\geq L\right\}\wedge 2T$
is larger than $T$ with probability greater than $\varkappa\in(0,1)$ given by the statement of ˜1.1. Notice that is always possible to find such $L$ since $H^{2+N_{1}}\subset C^{N_{1}}$ by Sobolev embedding. In this way, we have a $\mathbb{P}$-almost sure control over the size of $\|z_{t}\|_{N_{1}}$ on the time interval $t\in[0,\mathfrak{t}]$. However, since $z$ is temporally rough, we need to introduce a time mollification
(2.11) $z_{q}(t):=\int_{\mathbb{R}}r_{i_{q}}(t-s)z(s)\,ds,\quad t\in\mathbb{R}$
where z(s):=z(0)z(-s):=z(0) for all $s>0$, $r$ is a smooth time mollifier with support contained in $[0,1]$, and the parameter $i_{q}$ is defined as
(2.12) $i_{q}:=\begin{cases}\lambda_{0}^{m_{5}}\lambda_{1}^{m_{6}},\,&q=0,\\ \lambda_{q}^{m_{1}}\lambda_{q+1}^{m_{2}}\delta_{q}^{m_{3}}\delta_{q+1}^{m_{4}},\,&q\geq 1,\end{cases}$
where $(m_{1},m_{2},m_{3},m_{4})=\left(1/2-\delta\right)^{-1}\left(0,0,-2,7/2\right)$ and $(m_{5},m_{6})=(7\alpha/(1-2\delta),-m_{5}).$ The choice of the asymmetric mollifier guarantees that $z_{q}$ remains adapted to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$, extended to negative times identically equal to $\mathbb{F}_{0}$. Hereafter, we will always adopt this convention when referring to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$.
2.4. Induction scheme
The main idea of [Is22] is to obtain a solution of Euler equations satisfying the local energy inequality as a limit of smooth solutions of approximating systems, which include a Reynolds stress tensor $R$ in the momentum equation and a flux current $\varphi$ in the local energy inequality, and boil down to the Euler system when $R=0$ and $\phi=0$. In this sense, the size of the errors $R$ and $\phi$ give an indication of “how far” a solution of the approximating system is from being an exact solution of the Euler system.
In this work, the approximating system to (2.5) is slightly more involved, since we need to keep track of the random process $z$ affecting the dynamics. Throughout this work we shall deal with probabilistically strong solutions, i.e. our objects will be defined on the same stochastic basis $(\Omega,\mathbb{F},(\mathbb{F}_{t}),\mathbb{P})$ supporting the Wiener process $z$.
Definition 2.1.
A tuple of $\{\mathbb{F}_{t}\}$-progressively measurable smooth tensors $(v,p,R,\phi,z)$ is a dissipative Euler-Reynolds flow, with energy loss $E\in C^{1}_{b}$, if it satisfies the following system
(2.13) $\begin{cases}\partial_{t}v+\mathord{{\rm div}}\big((v+z)\otimes(v+z)\big)+\nabla p=\mathord{{\rm div}}(R),\\ \mathord{{\rm div}}(v)=0,\\ \partial_{t}\frac{|v|^{2}}{2}+v\cdot\nabla p+v\cdot\mathord{{\rm div}}\big((v+z)\otimes(v+z)\big)=-E^{\prime}+\frac{1}{2}D_{t}(\operatorname{Tr}(R))+\mathord{{\rm div}}(Rv)+R:\nabla z^{T}+\mathord{{\rm div}}(\phi),\end{cases}$
where $D_{t}:=\partial_{t}+(v+z)\cdot\nabla$ denotes the advective derivative.
We point out that our approximating Euler-Reynolds flow will solve (2.13) (in an analytically strong sense) also for negative times $t<0$, as we work with one-sided time mollifications in order to preserve adaptedness to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$. We shall impose zero external noise for negative times, so that the process $(v,p,R,\phi,z)$ will be deterministic for negative times (and thus progressively measurable with respect to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$).
As usual in convex integration schemes, we set up an iterative procedure to modify a given dissipative Euler-Reynolds flow $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$, $q\in\mathbb{N}$, in such a way to identify a new dissipative Euler-Reynolds flow $(v_{q+1},p_{q+1},R_{q+1},\phi_{q+1},z_{q+1})$ whose errors satisfy $\|R_{q+1}\|_{0}\ll\|R_{q}\|_{0}$ and $\|\phi_{q+1}\|_{0}\ll\|\phi_{q}\|_{0}$.
Lemma 2.2 (Iteration Lemma).
Let $\alpha\in(0,1/7)$ and $E\in C^{1}_{b}((-\infty,2T],\mathbb{R})$. Then there exist parameters $a,b$ as above, $C_{v},\overline{M}\in(1,\infty)$ and $n_{0}\in\mathbb{N}$ such that the following holds.
Assume that $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$, $q\in\mathbb{N}$, is a $\{\mathbb{F}_{t}\}$-progressively measurable dissipative Euler-Reynolds flow solving (2.13) on the time interval $(-\infty,\mathfrak{t}]$ and such that $\mathbb{P}$-almost surely the following estimates hold for all $t\leq\mathfrak{t}$:
(2.14) $\|v_{q}(t)\|_{N}$ $\leq C_{v}\overline{M}\lambda_{q}^{N}\delta_{q}^{1/2},$ $N\in\{1,2,\dots,n_{0}+3\},$
(2.15) $\|p_{q}(t)\|_{N}$ $\leq\overline{M}\lambda_{q}^{N}\delta_{q},$ $\|D_{t,q}p_{q}(t)\|_{N-2}\leq\overline{M}\lambda_{q}^{N-1}\delta_{q}^{3/2},$ $N\in\{1,2,\dots,n_{0}+3\},$
(2.16) $\|R_{q}(t)\|_{N}$ $\leq\overline{M}\lambda_{q}^{N-\gamma}\delta_{q+1},$ $\|D_{t,q}R_{q}(t)\|_{N-2}\leq\overline{M}\lambda_{q}^{N-1-\gamma}\delta_{q+1}\delta_{q}^{1/2},$ $N\in\{0,1,\dots,n_{0}+3\},$
(2.17) $\|\phi_{q}(t)\|_{N}$ $\leq\overline{M}\lambda_{q}^{N-3\gamma/2}\delta_{q+1}^{\frac{3}{2}},$ $\|D_{t,q}\phi_{q}(t)\|_{N-1}\leq\overline{M}\lambda_{q}^{N-3\gamma/2}\delta_{q+1}^{\frac{3}{2}}\delta_{q}^{1/2},$ $N\in\{0,1,2\},$
where $\gamma:=(b-1)^{2}$ and $D_{t,q}:=\partial_{t}+(v_{q}+z_{q})\cdot\nabla$.
Then, there exists a new tuple $(v_{q+1},p_{q+1},R_{q+1},\phi_{q+1},z_{q+1})$ of $\{\mathbb{F}_{t}\}$-progressively measurable tensors which solves (2.13), such that the estimates (2.14), (2.15), (2.16), (2.17) hold with $q$ replaced by $q+1$ and, in addition, $\mathbb{P}$-almost surely for every $t\leq\mathfrak{t}$:
(2.18) $\|v_{q+1}-v_{q}\|_{0}+\frac{1}{\lambda_{q+1}}\|v_{q+1}-v_{q}\|_{1}$ $\leq C_{v}\overline{M}\delta_{q+1}^{1/2}\lambda_{q}^{-\gamma/2},$
(2.19) $\|p_{q+1}-p_{q}\|_{0}+\frac{1}{\lambda_{q+1}}\|p_{q+1}-p_{q}\|_{1}$ $\leq\overline{M}\delta_{q+1}.$
We point out that, by definition of $\|\cdot\|_{-1}$ and $\|\cdot\|_{-2}$, estimates (2.15), (2.16) and (2.17) on the advective derivatives are void in these specific cases. Notice also that $\mathfrak{t}\leq 2T$ $\mathbb{P}$-almost surely by definition, so that the system (2.13) is meaningful on the time interval $(-\infty,\mathfrak{t}]$.
Since $\delta_{q}\to 0$ exponentially fast, the estimates above imply $\|R_{q}\|_{0},\|\phi_{q}\|_{0}\to 0$ as $q\to\infty$. Moreover, (2.18) and (2.19) on the increments and interpolation guarantee convergence $v_{q}\to v$ and $p_{q}\to p$ in a space of continuous-in-time, Hölder-in-space functions. In fact, given the previous lemma, we can prove ˜1.1 arguing as follows.
Proof of ˜1.1.
Step 1: Initialization. Let $E\in C^{1}_{loc}([0,\infty),\mathbb{R})$ be as in the statement of the theorem. Then we can find an extension of $E$ to negative times, still denoted $E$ with a little abuse of notation, such that $C^{1}_{b}((-\infty,2T],\mathbb{R})$. Consider the following dissipative Euler-Reynolds flow at step $q=0$:
(2.20) $(z_{0},v_{0},p_{0},R_{0},\phi_{0}):=\left(z_{0},0,\frac{|z_{0}|^{2}}{3},z_{0}\mathring{\otimes}z_{0}+\frac{2E}{3}\mathrm{Id},-\frac{|z_{0}|^{2}}{2}z_{0}\right),$
[…]
(2.21) $\|R_{0}(t)\|_{N}$ $\leq L^{2}+|E(t)|\leq\lambda_{0}^{N-\gamma}(L^{2}+|E(t)|)\lambda_{0}^{\gamma},$
(2.22) $\|D_{t,0}R_{0}(t)\|_{N-2}$ $\leq i_{0}^{-1/2-\delta}L^{2}+L^{3}+|E^{\prime}(t)|\leq\lambda_{0}^{N-1-\gamma}(L^{3}+|E^{\prime}(t)|)\lambda_{0}^{3\gamma/2},$
where we have used that $i_{0}^{-1/2-\delta}\leq\lambda_{0}^{1+\gamma/2}$, as well as the bounds on the current
(2.23) $\|\phi_{0}(t)\|_{N}\leq L^{3}\leq L^{3}\lambda_{0}^{N-3\gamma/2}\lambda_{0}^{3\gamma/2},\quad\|D_{t,0}\phi_{0}(t)\|_{N-1}\leq i_{0}^{-1/2-\delta}L^{3}+L^{4}\leq L^{4}\lambda_{0}^{N-3\gamma/2}\lambda_{0}^{2\gamma},$
and the pressure
(2.24) $\|p_{0}(t)\|_{N}\leq L^{2}\leq L^{2}\lambda_{0}^{N},\quad\|D_{t,0}p_{0}(t)\|_{N-2}\leq i_{0}^{-1/2-\delta}L^{2}+L^{3}\leq L^{3}\lambda_{0}^{N-1}\lambda_{0}^{\gamma/2}.$
Then, recalling that by definition $\delta_{0}=\delta_{1}=\lambda_{0}^{\gamma}>1$, the estimates above imply the existence of a constant $\overline{M}\in(1,\infty)$, depending only on $L$ and $E$, such that (2.14), (2.15), (2.16), (2.17) hold true $\mathbb{P}$-almost surely on the time interval $(-\infty,\mathfrak{t}]$.
Step 2: Construction of $(u_{t\wedge\mathfrak{t}},p_{t\wedge\mathfrak{t}})$. We iteratively apply Lemma˜2.2, starting from the initial tuple $(z_{0},v_{0},p_{0},R_{0},\phi_{0})$ defined in Step 1. We obtain a sequence of dissipative Euler-Reynolds flows, $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$, and by (2.18) we have that $v_{q}$ converges $\mathbb{P}$-almost surely to some $v$ in $C_{t}^{0}C^{\beta}_{x}$ for any $\beta<\alpha$. Similarly, by (2.19), $p_{q}$ converges $\mathbb{P}$-almost surely to some $p$ in $C_{t}^{0}C^{2\beta}_{x}$. Moreover, by construction $z_{q}$ converges $\mathbb{P}$-almost surely to $z$ uniformly on the time interval $0,\mathfrak{t}]$, while $\|R_{q}\|_{0}$ and $\|\phi_{q}\|_{0}$ vanish by (2.16) and (2.17). Notice that the tuple $(v,p,z)$ is $\{\mathbb{F}_{t}\}$-progressively measurable since each $(v_{q},p_{q},z_{q})$ is so, by Lemma˜2.2. Thus, we have built a solution of the system (2.5) on the time interval $0,\mathfrak{t}]$. By the discussion in Section˜2.1, we have that the process $(u,p)$ with $u:=v+z$ solves (SE) and (LEE) on the same time interval, and has the regularity stated in ˜1.1.
Step 2: Continuation of $(u,p)$. The proof is similar to [LuLuZh25, Corollary 1.8]. Let $\tilde{W}_{t}:=W_{t+\mathfrak{t}}-W_{\mathfrak{t}}$ be a cylindrical Wiener process on the filtration $\tilde{\mathbb{F}}_{t}:=\sigma(\{\tilde{W}_{s}\}_{s\leq t},u_{\mathfrak{t}},p_{\mathfrak{t}})$. We intend to apply [LuLuZh25, Theorem 1.7] with initial condtion $u^{in}:=u_{\mathfrak{t}}\in C^{\beta}_{x}$ $\mathbb{P}$-almost surely. This produces at least two distinct solutions $(\tilde{u}^{1},\tilde{p}^{1})$ and $(\tilde{u}^{2},\tilde{p}^{2})$ to (SE) with noise $Q^{1/2}d\tilde{W}$, $\mathbb{P}$-almost surely of class $\tilde{u}^{i}\in C^{0}_{t}C^{\vartheta}_{x}$ with arbitrary $\vartheta<\beta$. The pressure fields $\tilde{p}^{i}$ can be recovered from $\tilde{u}^{i}$ given by the aforementioned theorem by solving the equations $\Delta\tilde{p}^{i}=\mathord{{\rm div}}\mathord{{\rm div}}(\tilde{u}^{i}\otimes\tilde{u}^{i})$ subject to the boundary conditions $\int_{\mathbb{T}^{3}}\tilde{p}^{i}(t,x)dx=\int_{\mathbb{T}^{3}}p_{\mathfrak{t}}(x)dx$ for every $t\geq 0$, and are of class $\tilde{p}^{i}\in C_{t}C^{2\vartheta}_{x}$ $\mathbb{P}$-almost surely by [CoDR20, Corollary 3.3].
Then the glued processes $(u^{i}_{t},p^{i}_{t}):=(u_{t},p_{t})\mathbf{1}_{\{t\leq\mathfrak{t}\}}+(\tilde{u}^{i}_{t-\mathfrak{t}},\tilde{p}^{i}_{t-\mathfrak{t}})\mathbf{1}_{\{t>\mathfrak{t}\}}$, $i=1,2$, are solutions to (SE) that satisfy (LEE) by construction on the time interval $t\in[0,\mathfrak{t}]$. Since $\vartheta<\beta<\alpha$ is arbitrary, each $(u^{i},p^{i})$ has the space regularity stated in ˜1.1, concluding the proof.
Remark 2.3.
Let us stress that (2.14) and (2.18) imply that, for any $q\in\mathbb{N}$ and $N\leq n_{0}+3$,
(2.25) $\|u_{q}(t)\|_{N}+\|v_{q}(t)\|_{N}\lesssim\begin{cases}\lambda_{q}^{N}\delta_{q}^{1/2},\quad&N>0\\ 1\,,\quad&N=0\end{cases},$
where the implicit constant does not depend on the parameter $a$. This independence of $a$ will be crucial in the following sections. Indeed, we will be able to absorb various implicit constants emerging from the convex integration scheme by picking $a$ large enough. Since this remark is not trivial just for $N=0$, let us focus on that case. Indeed, since $a>2$ it holds for every $q\in\mathbb{N}$
(2.26) $\|v_{q}(t)\|_{0}$ $\leq\sum_{j=1}^{q}\|v_{j}(t)-v_{j-1}(t)\|_{0}\leq C_{v}\overline{M}\sum_{j=1}^{q}\delta_{j}^{1/2}\lambda_{j-1}^{-\gamma/2}$
$\leq C_{v}\overline{M}\sum_{j=1}^{q}\left(\frac{\lambda_{1}}{\lambda_{j}}\right)^{\alpha}\left(\frac{\lambda_{0}}{\lambda_{j-1}}\right)^{\gamma/2}\leq C_{v}\overline{M}\left(1+\sum_{j=2}^{q}a^{\alpha(b-b^{j})}\right)\leq C_{v}\overline{M}\left(1+\sum_{j=2}^{q}2^{\alpha(b-b^{j})}\right).$
3. Construction of the velocity perturbation
This section explains the construction of the perturbation $w:=v_{q+1}-v_{q}$, where $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$ is a solution to (2.13). In particular, $w$ is obtained by adapting the construction introduced in [DLK23] to our stochastic framework.
3.1. Mikado flows
The perturbation $w$ is built by combining highly oscillating vector fields modulated by scalar amplitudes, that will be chosen to cancel out $R_{q}$ and $\phi_{q}$ through the following two geometric lemmas.
Lemma 3.1.
Let $\mathcal{S}^{3}$ denote the space of $3\times 3$ symmetric matrices, and consider a family of vectors $\mathcal{F_{R}}:=\{k_{i}\}_{i=1}^{6}\subset\mathbb{Z}^{3}$ and a constant $C>0$ such that
(3.1) $\sum\limits_{i=1}^{6}k_{i}\otimes k_{i}=C\mathrm{Id},\quad\text{and }\{k_{i}\otimes k_{i}\}\text{ is a basis of }\mathcal{S}^{3}.$
Then, there exists a positive constants $N_{0}=N_{0}(\mathcal{F_{R}})$ such that for any $N\leq N_{0}$ we can find functions $\{\Gamma_{k_{i}}\}_{i=1}^{6}\subset\mathscr{C}^{\infty}(S_{N},(0,\infty))$, where $S_{N}:=\{K\in\mathcal{S}^{3}:|K-Id|_{\infty}\leq N\}$, such that for every $K\in S_{N}$ it holds
(3.2) $K=\sum\limits_{i=1}^{6}\Gamma_{k_{i}}^{2}(K)k_{i}\otimes k_{i}.$
Lemma 3.2.
Suppose that $\mathcal{F_{\phi}}:=\{k_{1},k_{2},k_{3}\}\subset\mathbb{Z}^{3}$ is an orthogonal frame and define
(3.3) $k_{4}:=-\sum\limits_{i=1}^{3}k_{i}.$
Then, for any $N_{0}>0$, there exist real-valued affine functions $\{\Lambda_{k_{i}}\}_{i=1}^{4}$ on $\mathbb{R}^{3}$ such that their restrictions to $\mathcal{V}_{N_{0}}:=\{u\in\mathbb{R}^{3}:|u|\leq N_{0}\}$ are of class $\mathscr{C}^{\infty}(\mathcal{V}_{N_{0}},[N_{0},\infty))$ and for every $u\in\mathcal{V}_{N_{0}}$ it holds
(3.4) $u=\sum\limits_{i=1}^{4}\Lambda_{k_{i}}(u)k_{i}.$
As in [DLK23], these two lemmas allow to find 27 disjoint families of directions $\{\mathcal{F}^{j}\}_{j\in\mathbb{Z}^{3}_{3}}$, indexed by equivalence classes $j\in\mathbb{Z}^{3}_{3}$ (triples of integers modulo $3$), such that any $\mathcal{F}^{j}$ can be decomposed as $\mathcal{F}^{j}=\mathcal{F}^{j}_{R}\cup\mathcal{F}^{j}_{\phi}$, where $\mathcal{F}^{j}_{R}$ and $\mathcal{F}^{j}_{\phi}$ satisfy the assumptions of Lemmas˜3.1 and 3.2, respectively. For any direction $f\in\mathcal{F}:=\bigcup_{j\in\mathbb{Z}^{3}_{3}}\mathcal{F}^{j}$ let us define
(3.5) $l_{f}:=\{x\in\mathbb{T}^{3}:x-\sigma f\in 2\pi\mathbb{Z}^{3}\text{ for some }\sigma\in\mathbb{R}\},$
which is nothing but the the line in direction $f$ passing through the origin, but seen as subset of $\mathbb{T}^{3}$. Moreover, le $\eta$ be a positive constant and $\psi_{f}\in C^{\infty}(\mathbb{T}^{3},\mathbb{R})$ be a function such that:
(3.6) $\operatorname{Supp}(\psi_{f})\subset B\left(l_{f},\frac{\eta}{10}\right),\quad f\cdot\nabla\psi_{f}=0,$
namely with support concentrated in the pipe of radius $\eta/10$ around the line $l_{f}$, and invariant by translations in the direction $f$. We assume the following renormalization properties:
(3.7) $\int_{\mathbb{T}^{3}}\psi_{f}dx=\int_{\mathbb{T}^{3}}\psi_{f}^{3}dx=0,\quad\fint_{\mathbb{T}^{3}}\psi_{f}^{2}dx=1,$ $\text{ for all }f\in\mathcal{F}_{R};$
(3.8) $\int_{\mathbb{T}^{3}}\psi_{f}dx=0,\quad\fint_{\mathbb{T}^{3}}\psi^{3}_{f}dx=1,$ $\text{ for all }f\in\mathcal{F}_{\phi}.$
3.2. Partition of unity and shifts
In the following we denote $Q(x,r):=\{y\in\mathbb{T}^{3}:|y-x|_{\infty}\leq r\}$ the hypercube in $\mathbb{T}^{3}$ of points whose box distance from $x$ is at most $r$. Let $\chi_{0}:\mathbb{T}^{3}\to\mathbb{R}_{+}$ and $\theta_{0}:\mathbb{R}\to\mathbb{R}_{+}$ be nonnegative smooth functions and suppose $\operatorname{Supp}(\chi_{0})\subset Q\left(0,9\pi/8\right)$ with $\chi_{0}\equiv 1\,\mbox{on}\,Q\left(0,7\pi/8\right)$, and similarly $\operatorname{Supp}(\theta_{0})\subset[-1/8,9/8]$ with $\theta_{0}\equiv 1\,\mbox{on}\,[1/8,7/8]$.
For every $n\in\mathbb{Z}^{3}$ and $m\in\mathbb{Z}$, define $\chi_{n}(x):=\chi_{0}(x-2\pi n)$ and $\theta_{m}(t)=\theta_{0}(t-m)$. The functions $\chi_{0},\theta_{0}$ can be chosen in such a way that
(3.9) $\sum_{n\in\mathbb{Z}^{3}}\chi_{n}^{6}(x)=1,\quad\forall x\in\mathbb{T}^{3};\qquad\sum_{m\in\mathbb{Z}}\theta_{m}^{6}(t)=1,\quad\forall t\in\mathbb{R}.$
Next, let $\nu:=\frac{1-7\alpha}{3\alpha}$ such that $(7+\nu)\alpha<1$ and $(4+2\nu)\alpha-1<-3\alpha$. Then, we introduce the parameters
(3.10) $\epsilon:=\begin{cases}C_{\epsilon}\lambda_{0}^{h_{5}}\lambda_{1}^{h_{6}},\,&q=0,\\ C_{\epsilon}\lambda_{q}^{h_{1}}\lambda_{q+1}^{h_{2}}\delta_{q}^{h_{3}}\delta_{q+1}^{h_{4}},\,&q\geq 1,\end{cases}\qquad C_{\epsilon}:=(C_{v}\overline{M}+L)^{-1}\min\left(\frac{1}{2},\frac{\inf_{j}N_{0}(\mathcal{F}_{R}^{j})}{6}\right),$
which represents the size of the temporale partition, and
(3.11) $\mu:=\begin{cases}\lambda_{0}^{u_{5}}\lambda_{1}^{u_{6}},\,&q=0\\ \lambda_{q}^{u_{1}}\lambda_{q+1}^{u_{2}}\delta_{q}^{u_{3}}\delta_{q+1}^{u_{4}},\,&q\geq 1\end{cases},$
which represents instead the size of the spatial partition, where the exponents $h_{i},u_{i}$ are given by:
(3.12) $h_{1}=u_{1}=-1,\quad h_{2}=u_{2}=0,\quad h_{3}=u_{3}=-(2+\nu),\quad u_{4}=-u_{3},\quad h_{4}=\frac{3+2\nu}{2},$
$h_{5}=u_{5}=-1/2,\quad h_{6}=u_{6}=-1/2.$
Indeed, given index sets
(3.13) $\mathcal{I}$ $=\{(m,n,f)\in\mathbb{Z}\times\mathbb{Z}^{3}\times\mathcal{F}:f\in\mathcal{F}^{n}\},$
$\mathcal{I}_{R}$ $=\{I\in\mathcal{I}:f\in\mathcal{F}_{R}^{n}\},$
$\mathcal{I}_{\phi}$ $=\{I\in\mathcal{I}:f\in\mathcal{F}_{\phi}^{n}\},$
then let us introduce
(3.14) $\chi_{I}(x):=\begin{cases}\chi_{n}^{2}(\frac{x}{\mu}),\quad I\in\mathcal{I}_{\phi},\\ \chi_{n}^{3}(\frac{x}{\mu}),\quad I\in\mathcal{I}_{R},\end{cases}\quad\mbox{ and }\quad\theta_{I}(t):=\begin{cases}\theta_{m}^{2}(\frac{t}{\epsilon}),\quad I\in\mathcal{I}_{\phi},\\ \theta_{m}^{3}(\frac{t}{\epsilon}),\quad I\in\mathcal{I}_{R}.\end{cases}$
Given a radial smooth function $\mathfrak{m}:\mathbb{R}^{3}\to\mathbb{R}$ supported in $B(0,2)$ and such that $\mathfrak{m}\equiv 1$ on $\overline{B(0,1)}$, consider the standard Littlewood-Paley operators acting on Fourier transforms as follows:
(3.15) $\mathcal{F}[\mathbf{P}_{\leq 2^{j}}f](\xi):=\mathfrak{m}(\xi\,2^{-j})\mathcal{F}[f](\xi),\quad\mathcal{F}[\mathbf{P}_{>2^{j}}f](\xi):=[1-\mathfrak{m}(\xi\,2^{-j})]\mathcal{F}[f](\xi),$
The symbol $\mathcal{F}$ here indicated the Fourier transform and should not be confused with the set of directions introduced in Section˜3.1. Given the vector $(e_{1},e_{2},e_{3},e_{4}):=(-1,0,-\frac{4+\nu}{2},\frac{4+\nu}{2}),(e_{5},e_{6}):=(-1/2,-1/2)$, let us introduce the frequency cutoff parameter
(3.16) $l:=\begin{cases}\lambda_{0}^{e_{5}}\lambda_{1}^{e_{6}},\,&q=0\\ \lambda_{q}^{e_{1}}\lambda_{q+1}^{e_{2}}\delta_{q}^{e_{3}}\delta_{q+1}^{e_{4}},\quad&q\geq 1\\ \end{cases},$
and denote
(3.17) $\mathbf{P}_{\leq l^{-1}}:=\mathbf{P}_{\leq 2^{S}},\quad S:=\sup\{j\in\mathbb{Z}\,:\,2^{j}\leq l^{-1}\}.$
Incidentally, we also define here the Littlewood-Paley operator projecting on the frequency shell of size approximately equal to $2^{j}$, namely:
(3.18) $\mathbf{P}_{2^{j}}f:=\mathbf{P}_{>2^{j-1}}-\mathbf{P}_{>2^{j}},$
that will be useful later. Then we define the spatial regularization of our velocity fields in the following way:
(3.19) $v_{l}:=\mathbf{P}_{\leq l^{-1}}[v_{q}],\quad z_{l}:=\mathbf{P}_{\leq l^{-1}}[z_{q}],\quad u_{l}:=v_{l}+z_{l}.$
Denote $D_{t,l}:=\partial_{t}+u_{l}\cdot\nabla$ the advective derivative with respect to the velocity field $u_{l}$. Finally, for any $I\in\mathcal{I}_{m}:=\{(m,n,f):n\in\mathbb{Z}^{3},f\in\mathcal{F}^{n}\}$, let us consider the flow $\xi_{I}$ solving
(3.20) $\begin{cases}D_{t,l}\xi_{I}(t,x)=0,\\ \xi_{I}(\epsilon(m-1/8),x)=x.\end{cases}$
3.3. Perturbation of the velocity
The principal part of the perturbation $w=v_{q+1}-v_{q}$ is defined as follows
(3.21) $w_{o}(t,x):=\sum_{I\in\mathcal{I}}\theta_{I}(t)\chi_{I}(\xi_{I}(t,x))a_{I}(t,x)\nabla\xi_{I}(t,x)^{-1}f_{I}\psi_{I}(\lambda_{q+1}\xi_{I}(t,x))=:\sum_{I\in\mathcal{I}}w_{I}^{o}(t,x),$
where the index $o$ stands for oscillatory. The amplitudes $a_{I}$ will be defined in the following Section˜3.4, and $\psi_{I}$ is given by
(3.22) $\psi_{I}(x)=\psi_{f_{I}}(x-s_{I}),$
where $f_{I}$ is the direction appearing in the index $I$ (i.e. $I=(m,n,f_{I})$ for some $m,n$), and the set of shifts $\{s_{I}\}_{I\in\mathcal{I}}\subset\mathbb{R}^{3}$ is chosen such that the following conditions hold:
  • $\operatorname{Supp}w_{I}^{o}\cap\operatorname{Supp}w_{J}^{o}=\emptyset$ for all $I\neq J$;
  • For every $I=(m,n,f)$, the shift $s_{I}$ satisfies $s_{I}=s_{m,n}+p_{n,f}$ for some $s_{m,n}$ depending only on $m,n\in\mathbb{Z}\times\mathbb{Z}^{3}$ and $p_{n,f}$ depending only on $f\in\mathcal{F}^{n}$;
  • $s_{m,n}=s_{m,n^{\prime}}$ for $(n^{\prime}-n)\mu\in\mathbb{Z}^{3}$.
The proof of the existence of such a family of shifts is a simple adaptation of [DLK23, Proposition 3.5], with only minor modifications needed to take the randomness of $u_{l}$ into account, and is omitted for the sake of brevity. As in [DLK23], we can expand $\psi_{I},\psi^{2}_{I},\psi^{3}_{I}$ in Fourier series as follows:
(3.23) $\psi_{I}=\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{b}_{I,k}e^{ik\cdot x},\quad\psi^{2}_{I}=\mathring{c}_{I,0}+\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{c}_{I,k}e^{ik\cdot x},\quad\psi^{3}_{I}=\mathring{d}_{I,0}+\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{d}_{I,k}e^{ik\cdot x},$
where (3.6) implies that $\mathring{b}_{I,k}(f_{I}\cdot k)=\mathring{c}_{I,k}(f_{I}\cdot k)=\mathring{d}_{I,k}(f_{I}\cdot k)=0$. Then:
(3.24) $w_{o}$ $=\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}b_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
(3.25) $w_{o}\otimes w_{o}$ $=c_{0}+\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}c_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
(3.26) $\frac{|w_{o}|^{2}}{2}w_{o}$ $=d_{0}+\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}d_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
where the coefficients $b_{m,k},c_{m,k},d_{m,k}$ are given by
(3.27) $b_{m,k}=\theta_{m}(t)\sum_{I\in\mathcal{I}_{m}}\chi_{I}(\xi_{I}(t,x))a_{I}(x,t)\overline{f_{I}}(x,t)\,\mathring{b}_{I,k},$
(3.28) $c_{m,k}=\theta^{2}_{m}(t)\sum_{I\in\mathcal{I}_{m}}\chi^{2}_{I}(\xi_{I}(x,t))a_{I}^{2}(t,x)\left(\overline{f_{I}}\otimes\overline{f_{I}}\right)(t,x)\,\mathring{c}_{m,k},$
(3.29) $d_{m,k}=\theta^{3}_{m}(t)\sum_{I\in\mathcal{I}_{m}}\chi^{3}_{I}(\xi_{I}(x,t))a_{I}^{3}(t,x)\left(\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}\right)(t,x)\,\mathring{d}_{m,k},$
with
(3.30) $\overline{f_{I}}(x,t):=(\nabla\xi_{m})^{-1}(x,t)f_{I},$
whereas the zero modes $c_{0},d_{0}$ are given by
(3.31) $c_{0}=\sum_{I\in\mathcal{I}}\theta^{2}_{I}(t)\chi_{I}^{2}(\xi(x,t))a_{I}^{2}(t,x)\left(\overline{f_{I}}\otimes\overline{f_{I}}\right)(t,x)\fint_{\mathbb{T}^{3}}\psi_{I}(t,y)^{2}dy,$
(3.32) $d_{0}=\sum_{I\in\mathcal{I}}\theta^{3}_{I}(t)\chi_{I}^{3}(\xi(x,t))a_{I}^{3}(t,x)\left(\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}\right)(t,x)\fint_{\mathbb{T}^{3}}\psi_{I}(t,y)^{3}dy.$
At last, we need to define a compressibility corrector term $w_{c}$ such that $\text{div}(w_{c}+w_{o})=0$. Let
(3.33) $w_{c}:=$ $\frac{1}{\lambda_{q+1}}\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}e_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
(3.34) $e_{m,k}=$ $\sum_{I\in\mathcal{I}_{m}}\nabla\left(\theta_{m}(t)\chi_{I}(\xi_{I}(t,x))a_{I}(t,x)\mathring{b}_{I,k}\right)\times\left(\nabla\xi_{I}^{T}(t,x)\frac{ik\times f_{I}}{|k|^{2}}\right).$
3.4. Choice of the amplitudes
The amplitudes $a_{I}$ are built differently based on $I$ being in $\mathcal{I}_{R}$ or in $\mathcal{I}_{\phi}$. First of all, consider a temporal mollifier $\eta$ with support in [1,0][-1,0], and a time mollification parameter
(3.35) $l_{\rm temp}:=\begin{cases}\lambda_{0}^{s_{5}}\lambda_{1}^{s_{6}},\,&q=0\\ \lambda_{q}^{s_{1}}\lambda_{q+1}^{s_{2}}\delta_{q}^{s_{3}}\delta_{q+1}^{s_{4}},\,&q\geq 1\end{cases},$
with $(s_{1},s_{2},s_{3},s_{4})=(-1,0,-\frac{4+\nu}{2},\frac{3+\nu}{2}),\,(s_{5},s_{6})=(-1/2,-1/2)$. The asymmetric choice for the support of $\eta$ guarantees that convolutions with $\eta$ preserve $\{\mathbb{F}_{t}\}_{t\geq 0}$ adaptedness. Then, define
(3.36) $R_{l}(t,x)$ $=\int_{\mathbb{R}}(\mathbf{P}_{\leq l^{-1}}R_{q})\left(t+s,\mu(t;t+s,x)\right)\eta_{l_{\rm temp}}(s)ds,$
(3.37) $\phi_{l}(t,x)$ $=\int_{\mathbb{R}}(\mathbf{P}_{\leq l^{-1}}\phi_{q})\left(t+s,\mu(t;t+s,x)\right)\eta_{l_{\rm temp}}(s)ds,$
where $\mu(t;\cdot,\cdot)$ is the unique solution to the transport equation
(3.38) $\begin{cases}\partial_{r}\mu(t;r,x)=u_{l}(r,\mu(t;r,x)),\quad r\geq t,\\ \mu(t;t,x)=x.\end{cases}$
Notice that this mollification along the Lagrangian flow implies that
(3.39) $D_{t,l}R_{l}(t,x)$ $=-\int_{\mathbb{R}}(\mathbf{P}_{\leq l^{-1}}R_{q})\left(t+s,\mu(t;t+s,x)\right)\eta_{l_{\rm temp}}^{\prime}(s)ds,$
$D_{t,l}\phi_{l}(t,x)$ $=-\int_{\mathbb{R}}(\mathbf{P}_{\leq l^{-1}}\phi_{q})\left(t+s,\mu(t;t+s,x)\right)\eta_{l_{\rm temp}}^{\prime}(s)ds.$
First case: $I\in\mathcal{I}_{\phi}$. We define $a_{I}$ to cancel out $\phi_{l}$ through $d_{0}$. Indeed, consider
(3.40) $a_{I}(t,x):=|\overline{f_{I}}|^{-2/3}2^{1/3}\Lambda_{f_{I}}^{1/3}\left(-\frac{\nabla\xi_{I}(t,x)\phi_{l}(t,x)}{\overline{M}^{3}\lambda_{q}^{-3\gamma/2}\delta_{q+1}^{3/2}}\right)\overline{M}(\lambda_{q}^{-3\gamma/2}\delta_{q+1}^{3/2})^{1/3}.$
Notice that $\epsilon\|u_{l}(t)\|_{1}<1/2$ implies $|\overline{f}_{I}|>1/2$ by (3.30). Then, (3.8) and (3.32) plus the disjoint supports property gives
(3.41) $d_{0}$ $=\sum_{I\in\mathcal{I_{\phi}}}\theta^{3}_{I}(t)\chi_{I}^{3}(\xi(x,t))a_{I}^{3}(t,x)\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}(t,x)$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))\nabla\xi_{m}^{-1}(t,x)\sum_{\begin{subarray}{c}I\in\mathcal{I}_{\phi}\\ I=(m,n,f_{I})\end{subarray}}\overline{M}^{3}\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}\Lambda_{f_{I}}\left(-\frac{\nabla\xi_{I}(t,x)\phi_{l}(t,x)}{\overline{M}^{3}\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}}\right)f_{I}$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))(-\phi_{l}(t,x))=-\phi_{l}(t,x),$
where the second-to-last equality is due to Lemma˜3.2 applied with $N_{0}=e$, while the last one is due to the space and time partitions of unity.
Second case: $I\in\mathcal{I}_{R}$. We define $a_{I}$ to cancel out $R_{l}$ through $c_{0}$. Indeed, let $I=(m,n,f)$ and define
(3.42) $a_{I}(t,x)$ $:=\Gamma_{f}\left(\mathrm{Id}+(\nabla\xi_{I}\nabla\xi_{I}^{T}-\mathrm{Id})-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\rho^{1/2},$
(3.43) $\quad\rho$ $:=\frac{C\,\overline{M}^{2}\lambda_{q}^{-\gamma}\delta_{q+1}}{\inf_{j}N_{0}(\mathcal{F}_{R}^{j})},$
where $\Gamma_{f}$ and $N_{0}(\mathcal{F}_{R}^{j})$ are given by Lemma˜3.1, $C>0$ is a constant depending on the $\Gamma_{f}$’s functions and $\sup_{f\in\mathcal{F}}|f|_{\infty}$, and
(3.44) $M_{m,n}:=\sum_{\begin{subarray}{c}J\in I_{\phi},\\ J=(m,n,f_{J}),\\ |m-m^{\prime}|_{\infty}\leq 1,\\ |n-n^{\prime}|_{\infty}\leq 1\end{subarray}}a_{J}^{2}(t,x)\chi_{J}^{2}(\xi_{J}(t,x))\theta_{J}^{2}(t)\left(\overline{f_{J}}\otimes\overline{f_{J}}\right)(t,x).$
Then, (3.7) and (3.25) imply that
(3.45) $c_{0}$ $=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))\left[\sum_{I\in\mathcal{I}_{m,n,R}}a_{I}^{2}(t,x)\nabla\xi_{I}^{-1}(t,x)(f_{I}\otimes f_{I})\nabla\xi_{I}^{-T}(t,x)+M_{m,n}\right]$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))\left[\rho\nabla\xi_{I}^{-1}\left(\nabla\xi_{I}\nabla\xi_{I}^{T}-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\nabla\xi_{I}^{-T}+M_{m,n}\right]$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))\left[\rho\mathrm{Id}-R_{l}\right]=\rho\mathrm{Id}-R_{l},$
where the second to last equality comes from Lemma˜3.1 under the additional assumption
(3.46) $\epsilon\|u_{l}(t)\|_{1}\leq\frac{\inf_{j}N_{0}(\mathcal{F}_{R}^{j})}{6}.$
3.5. Estimates on the velocity perturbation
With the construction above, we are able to control the size of the perturbation $w$, as stated in the following key result. Since its proof is highly technical and involved, the reader could skip it at a first reading without losing much insight on the remainder of this work. In brief, the main takeaway of this result is that taking the advective derivative $D_{t,l}$ effectively “costs” a factor $\epsilon^{-1}$, whereas a space derivative “costs” a factor $\lambda_{q+1}$.
Proposition 3.3.
For $s\in\{0,1,2\}$, $N\leq n_{0}+3$ and for every $t\leq\mathfrak{t}$, it holds
(3.47) $\epsilon^{s}\|D_{t,l}^{s}w_{o}(t)\|_{N}\lesssim\lambda_{q+1}^{N}\rho^{1/2},$
(3.48) $\epsilon^{s}\|D_{t,l}^{s}w_{c}(t)\|_{N}\lesssim(\mu\lambda_{q+1})^{-1}\lambda_{q+1}^{N}\rho^{1/2},$
and moreover there exists a finite constant $C_{2}$, independent on $C_{v}$, $\overline{M}$ and $L$, such that
(3.49) $\|w(t)\|_{N}\leq C_{2}\,\lambda_{q+1}^{N}\rho^{1/2}.$
In addition, for all $N\leq n_{0}+3$ and $t\leq\mathfrak{t}$, we have
(3.50) $\|u_{q+1}(t)-u_{q}(t)\|_{N}\lesssim\,\lambda_{q+1}^{N}\rho^{1/2},$
(3.51) $\|u_{q+1}(t)-u_{l}(t)\|_{N}\lesssim\,\lambda_{q+1}^{N}\rho^{1/2}.$
3.5.1. Proof of Proposition˜3.3: Preliminary estimates
We report some estimates that will be helpful in the proof of Proposition˜3.3, as well as in the remainder of the paper.
Lemma 3.4.
For any $N\leq n_{0}+4$ and $t\leq\mathfrak{t}$, we have
(3.52) $\|z_{l}(t)-z_{q+1}(t)\|_{N}\vee\|z_{q}(t)-z_{q+1}(t)\|_{N}\lesssim L\,i_{q}^{1/2-\delta};$
(3.53) $\|\partial_{t}z_{q}(t)\|_{N}\lesssim L\,i_{q}^{-1/2-\delta}.$
Proof.
The estimate for $\|z_{q}(t)-z_{q+1}(t)\|_{N}$ is standard, while the one for $\|z_{l}(t)-z_{q+1}(t)\|_{N}$ is due to
(3.54) $\|z_{l}(t)-z_{q}(t)\|_{N}\lesssim l\|z(t)\|_{N+1},\quad l\leq i_{q}^{1/2-\delta}.$
Instead, (3.53) follows from noticing that the time mollifier $\alpha$ used in (2.11) has compact support, implying
(3.55) $\int_{\mathbb{R}}r^{\prime}(s)z(t)ds=0,$
and giving in particular the bound
(3.56) $|\partial_{t}z_{q}(t)|$ $=\left|\partial_{t}z_{q}(t)-i_{q}^{-2}\int_{\mathbb{R}}r^{\prime}(s\,i_{q}^{-1})z(t)\,ds\right|$
(3.57) $=i_{q}^{-2}\left|\int_{\mathbb{R}}r^{\prime}(s\,i_{q}^{-1})[z(t-s)-z(t)]\,ds\right|$
(3.58) $=i_{q}^{-1}\left|\int_{\mathbb{R}}r^{\prime}(s)[z(t-i_{q}s)-z(t)]\,ds\right|\lesssim Li_{q}^{-1/2-\delta}.$
Lemma˜3.4 and (2.25) imply the following estimates, that will be useful later on.
Corollary 3.5.
For $N\leq n_{0}+3$ and $t\leq\mathfrak{t}$, we have that
(3.59) $\|D_{t,q+1}(z_{q}-z_{l})(t)\|_{N}$ $\lesssim l\left(i_{q}^{-1/2-\delta}+(\lambda_{q+1}^{N}\delta_{q+1}^{1/2}\vee 1)\right),$
(3.60) $\|D_{t,q+1}(z_{q+1}-z_{q})(t)\|_{N}$ $\lesssim i_{q+1}^{-1/2-\delta}+i_{q}^{1/2-\delta}(\lambda_{q+1}^{N}\delta_{q+1}^{1/2}\vee 1).$
Lemma 3.6.
For $N\leq n_{0}+1$ and $t\leq\mathfrak{t}$, we have
(3.61) $\|D_{t,l}(u_{q}-u_{l})(t)\|_{N}+\|D_{t,l}(v_{q}-v_{l})(t)\|_{N}\lesssim l^{1-N}\lambda_{q}^{2}\delta_{q}.$
Proof.
First of all, we have that
(3.62) $D_{t,q}u_{q}=\partial_{t}z_{q}+\mbox{div}(R_{q})+\nabla p_{q}.$
Therefore, the bounds (2.15), (2.16), [DLK23, Lemma A.3], $\lambda_{q}\delta_{q}^{1/2}\geq L$, $i_{q}^{-1/2-\delta}\leq\lambda_{q}^{2}\delta_{q}$ and $\lambda_{q}<l^{-1}$ imply
(3.63) $\|D_{t,l}(u_{q}-u_{l})\|_{N}$ $\leq\|\mathbf{P}_{>l^{-1}}D_{t,l}u_{q}\|_{N}+\|[\mathbf{P}_{>l^{-1}},u_{l}\cdot\nabla]u_{q}\|_{N}$
$\lesssim l\|D_{t,q}u_{q}\|_{N+1}+l\|(u_{l}-u_{q})\cdot\nabla u_{q}\|_{N+1}+l^{1-N}\|u_{q}\|_{1}^{2}\lesssim l^{1-N}\lambda_{q}^{2}\delta_{q}.$
The same computations and Lemma˜3.4 yield the result for $\|D_{t,l}(v_{q}-v_{l})(t)\|_{N}$. ∎
Lemma 3.7.
For any $N\leq n_{0}+3$ and $t\leq\mathfrak{t}$, we have
(3.64) $\|D_{t,l}\nabla v_{l}(t)\|_{N}+\|D_{t,l}\nabla u_{l}(t)\|_{N}\lesssim l^{-N}\lambda^{2}_{q}\delta_{q}.$
Proof.
Since it holds
(3.65) $D_{t,l}(\nabla u_{l})=\nabla(\partial_{t}u_{l}+u_{l}\cdot\nabla u_{l})-(\nabla u_{l})^{2}=\nabla\left(\mathord{{\rm div}}(R_{l}+R_{comm})-\nabla p_{l}+\partial_{t}z_{l}\right)-(\nabla u_{l})^{2},$
where $(\nabla u_{l})^{2}$ denotes the standard product between matrices $(\nabla u_{l})^{2}]^{i,j}=\sum_{k}\partial_{i}u^{k}\partial_{k}u^{j}$, then by [DLK23, Lemma A.3],Lemma˜3.4, and the bound $L\,i_{q}^{-1/2-\delta}\leq\lambda^{2}_{q}\delta_{q}$ we have:
(3.66) $\|D_{t,l}\nabla u_{l}(t)\|_{N}\lesssim l^{-N}\lambda^{2}_{q}\delta_{q}+L\,i_{q}^{-1/2-\delta}\lesssim l^{-N}\lambda^{2}_{q}\delta_{q}.$
Then, the estimate (3.64) on $\|D_{t,l}\nabla v_{l}(t)\|_{N}$ follows from the identity $D_{t,l}\nabla v_{l}(t)=D_{t,l}\nabla u_{l}(t)-D_{t,l}\nabla z_{l}(t)$ and computations similar to those above. ∎
3.5.2. Proof of Proposition˜3.3: Estimates on the oscillatory and compressible Fourier coefficients
Recall the definition of the parameters $\epsilon$ and $\mu$, given respectively in (3.10) and (3.11).
Let us start with estimates aimed to control the Fourier coefficients appearing $w$ and in the products $w_{o}\otimes w_{o}$ and $|w_{o}|^{2}w_{o}$, cf. Equations (3.24), (3.25), (3.26), (3.33).
The key takeaway of the following lemma is that taking the advective derivative of the building blocks costs a factor $\epsilon^{-1}$, whereas taking a space derivative costs a factor $\mu^{-1}$.
Lemma 3.8.
For $s\in\{0,1,2\}$ and $N\leq n_{0}+3$ , it holds for all $t\leq\mathfrak{t}$ that
(3.67) $\epsilon^{s}\|D_{t,l}^{s}b_{m,k}(t)\|_{N}\lesssim\sup_{I\in\mathcal{I}}|\mathring{b}_{I,k}|\rho^{1/2}\mu^{-N},$
(3.68) $\epsilon^{s}\|D_{t,l}^{s}c_{m,k}(t)\|_{N}\lesssim\sup_{I\in\mathcal{I}}|\mathring{c}_{I,k}|\rho\,\mu^{-N},$
(3.69) $\epsilon^{s}\|D_{t,l}^{s}d_{m,k}(t)\|_{N}\lesssim\sup_{I\in\mathcal{I}}|\mathring{d}_{I,k}|\rho^{3/2}\mu^{-N},$
(3.70) $\epsilon^{s}\|D_{t,l}^{s}e_{m,k}(t)\|_{N}\lesssim\sup_{I\in\mathcal{I}}|\mathring{b}_{I,k}|\rho^{1/2}\mu^{-N-1}.$
Proof.
Let us show the claim for $b_{m,k}$. The estimates on the other coefficients follow from similar inequalities. Due to the expression (3.27) and the spatial partition of unity, we can focus on a single addend corresponding to $I\in\mathcal{I}_{m}$ and $k\in\mathbb{Z}^{3}\setminus\{0\}$:
(3.71) $X_{I}(t,x):=\theta_{m}(t)\chi_{I}(\xi_{I}(t,x))a_{I}(x,t)\overline{f}_{I}(x,t)\mathring{b}_{I,k}.$
First, we have that
(3.72) $|D_{t,l}^{s}\theta_{m}(t)|\lesssim\epsilon^{-s}.$
Second, by Proposition˜A.1 and (3.64), together with the assumed bound $\epsilon\|u_{l}(t)\|_{1}<1$ and the algebraic identities
(3.73) $D_{t,l}(\nabla\xi_{I})^{-1}=\nabla u_{l}(\nabla\xi_{I})^{-1},\quad D_{t,l}^{2}(\nabla\xi_{I})^{-1}=D_{t,l}(\nabla u_{l})(\nabla\xi_{I})^{-1}+(\nabla u_{l})^{2}(\nabla\xi_{I})^{-1},$
imply that for every $K\in\{0,1\dots,N\}$ we have
(3.74) $\|D_{t,l}^{s}\overline{f_{I}}(t)\|_{K}\lesssim l^{-K}(\lambda_{q}\delta_{q}^{1/2})^{s},\quad\text{for }s\in\{0,1,2\}.$
Third, by Proposition˜A.1, (3.20) and [BD15+, Proposition C.1], we have
(3.75) $\|D_{t,l}^{s}\chi_{I}(\xi_{I})\|_{K}\lesssim\mu^{-K}\mathbf{1}_{\{s=0\}}.$
Now, we need to estimate $\|D_{t,l}^{s}a_{I}\|_{K}$ for $K\in\{0,1\dots,N\}$, either for $I\in\mathcal{I}_{\phi}$ or $I\in\mathcal{I}_{R}$. Before treating each case separately, let us recall some useful inequalities valid for general $C^{K}$ maps $g:\mathbb{T}^{3}\to\mathbb{R}^{k}$ and $\Lambda:\mathbb{R}^{k}\to\mathbb{R}^{n}$, for general $k,n$. First, it is easy to check that the advective derivative of $\Lambda(g):\mathbb{T}^{3}\to\mathbb{R}^{n}$ is given by $D_{t,l}\Lambda(g)=(\nabla\Lambda)(g)\cdot D_{t,l}g$. In particular, by Faà di Bruno’s formula we obtain for every $K\in\{0,1\dots,N\}$:
(3.76) $\|D_{t,l}\Lambda(g)\|_{K}\lesssim\sum_{K_{1}+K_{2}\leq K}\|D_{t,l}g\|_{K_{1}}\|(\nabla\Lambda)(g)\|_{K_{2}};$
(3.77) $\|D_{t,l}^{2}\Lambda(g)\|_{K}\lesssim\sum_{K_{1}+K_{2}\leq K}\|D_{t,l}^{2}g\|_{K_{1}}\|(\nabla\Lambda)(g)\|_{K_{2}}+\|D_{t,l}g\otimes D_{t,l}g\|_{K_{1}}\|(\nabla^{2}\Lambda)(g)\|_{K_{2}}.$
First case: $I\in\mathcal{I}_{\phi}$. Let $\rho_{1}:=\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}$. Recalling the definition of $\rho$ given in (3.43), one immediately has $\rho_{1}\lesssim\rho^{3/2}$ up to multiplicative constants independent of $\epsilon,\mu$. For $s=0$, [BD15+, Proposition C.1] implies that
(3.78) $\|a_{I}\|_{K}\lesssim l^{-K}\rho_{1}^{1/3}\lesssim l^{-K}\rho^{1/2}.$
For $s=1$ instead, we have by (3.73), Proposition˜A.1 and Eq.˜3.36:
(3.79) $\|D_{t,l}(\nabla\xi_{I}\phi_{l})\rho_{1}^{-1/3}\|_{K}\lesssim l_{\rm temp}^{-1}\,l^{-K}.$
On the other hand, [BD15+, Proposition C.1] implies that for $k\in\{1,2\}$
(3.80) $\left\|(\nabla^{k}\Lambda_{f_{I}}^{1/3})\left((\nabla\xi_{I}\phi_{l})\rho_{1}^{-1/3}\right)\right\|_{K}\lesssim l^{-K},$
where the implicit constant depends on $\|\nabla^{k}\Lambda_{f_{I}}^{1/3}\|_{K}$. Therefore, (3.76), (3.79), (3.80) imply that
(3.81) $\|D_{t,l}a_{I}\|_{K}$ $\lesssim\sum_{K_{1}+K_{2}\leq K}\rho_{1}^{1/3}\|D_{t,l}(\nabla\xi_{I}\phi_{l})\rho_{1}^{-1/3}\|_{K_{1}}\left\|(\nabla\Lambda_{f_{I}}^{1/3})\left((\nabla\xi_{I}\phi_{l})\rho_{1}^{-1/3}\right)\right\|_{K_{2}}$
$\lesssim\rho_{1}^{1/3}l_{\rm temp}^{-1}l^{-K}\lesssim\rho^{1/2}l_{\rm temp}^{-1}l^{-K}.$
Analogously for $s=2$ we have by (3.73) and (3.36)
(3.82) $\|D_{t,l}^{2}(\nabla\xi_{I}\phi_{j})\rho_{1}^{-1/3}\|_{K}\lesssim l_{\rm temp}^{-2}\,l^{-K}.$
Then, (3.77), (3.82) and (3.80) imply that
(3.83) $\|D_{t,l}^{2}a_{I}\|_{N}\lesssim\rho_{1}^{1/3}l_{\rm temp}^{-2}\,l^{-N}\lesssim\rho^{1/2}l_{\rm temp}^{-2}\,l^{-N}.$
To summarize, recalling the decomposition (3.71), we can combining all the estimates given in (3.72), (3.75), (3.74), (3.78), (3.81), and (3.83) above and finally obtain
(3.84) $\|D_{t,l}^{s}{X}_{I}\|_{N}\lesssim\epsilon^{-s}\sup_{I}|\mathring{b}_{I,k}|\rho^{1/2}\mu^{-N}.$
Second case: $I\in\mathcal{I}_{R}$. The desired estimates follow from the same set of inequalities used in the case $I\in\mathcal{I}_{\phi}$. The main difference, due to (3.42), is that we now need to estimate the quantity
(3.85) $\left\|D_{t,l}^{s}\left((\nabla\xi_{I}\nabla\xi_{I}^{T}-\mathrm{Id})-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\right\|_{N},$
where $M_{m,n}$ is given by Eq.˜3.44. Then, (3.73), (3.36), (3.84) imply that
(3.86) $\|D_{t,l}^{s}\left((\nabla\xi_{I}\nabla\xi_{I}^{T}-\mathrm{Id})-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\|_{N}\lesssim\epsilon^{-s}\mu^{-N}.$
Let us stress that we have exploited the bound (3.84) since $M_{m,n}(t,x)$ is given by a finite summation of $X_{J}$ for $J\in\mathcal{I}_{\phi}$. Therefore,
(3.87) $\|D_{t,l}^{s}X_{I}\|_{N}\lesssim\epsilon^{-s}\sup_{I}|\mathring{b}_{I,k}|\rho^{1/2}\mu^{-N}.$
Remark 3.9.
In particular, the choice of $C_{\epsilon}$ implies that there exists $C_{1}$ independent on $C_{v}$, $\overline{M}$ and $L$ such that
(3.88) $\|b_{m,k}(t)\|_{N}+\mu^{-1}\|e_{m,k}(t)\|_{N}\leq C_{1}\rho^{1/2}\mu^{-N}.$
We point out that the estimate above depends on $\epsilon$ because of the construction: whenever one has to bound $\xi_{m}$, terms of the form $\epsilon\|u_{q}\|_{1}l^{-j}$ appear; therefore the constant $C_{\epsilon}$ permits to compensate for the dependence of these terms on the parameters $C_{v}$ and $\overline{M}$ implicit in the quantity $\|v_{q}\|_{1}\leq C_{v}\overline{M}\lambda_{q}\delta_{q}^{1/2}$.
3.5.3. Proof of Proposition˜3.3: Conclusion
Let us start from the estimates on $w$. It is enough to show the result for $w_{o}$, as the claim for $w_{c}$ follows by similar arguments. Notice that
(3.89) $D_{t,l}^{s}w_{o}=\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}D_{t,l}^{s}b_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
where the sum over $m\in\mathbb{Z}$ is effectively a finite sum because the coefficients $b_{m,k}$ have disjoint supports for $|m-m^{\prime}|>1$. Furthermore, for every fixed $m\in\mathbb{Z}$ we have
(3.90) $\|e^{i\lambda_{q+1}\xi_{m}(t)\cdot k}\|_{N}\lesssim|k|^{N}\lambda^{N}_{q+1},$
and thus the desired inequality follows from Lemma˜3.8 and the analog of [DLK23, Equation (3.24)].
Moreover, the estimates (3.47), (3.48), (3.52), (2.14), $i_{q}^{1/2-\delta}\leq\rho^{1/2}$ and $l\lambda_{q}\delta_{q}^{1/2}\leq\rho^{1/2}$ imply the bounds (3.50) and (3.51) on $u$.
4. Construction of the new Reynolds stress and pressure
Based on the previous section, we have defined $v_{q+1}=v_{q}+w_{o}+w_{c}$. Now, we focus on identifying the new Reynolds stress $R_{q+1}$ and pressure $p_{q+1}$. Using $v_{q+1}=v_{q}+w_{o}+w_{c}$ and the equation $\partial_{t}v_{q}+\mathord{{\rm div}}\big(u_{q}\otimes u_{q}\big)+\nabla p_{q}=\mathord{{\rm div}}\big(R_{q}\big)$, we obtain that any putative $R_{q+1}$ and $p_{q+1}$ satisfying the momentum equation in (2.13) must necessarily satisfy also
(4.1) $-\nabla p_{q+1}+\mathord{{\rm div}}(R_{q+1})$ $=\partial_{t}v_{q+1}+\mathord{{\rm div}}\big(u_{q+1}\otimes u_{q+1}\big)$
$=\partial_{t}w+\mathord{{\rm div}}(w_{o}\otimes w_{o}+R_{l})-\nabla p_{q}+\mathord{{\rm div}}(w_{c}\otimes w+w_{o}\otimes w_{c})$
$\quad+u_{l}\cdot\nabla w+w\cdot\nabla u_{l}+\mathord{{\rm div}}\big((v_{q}-v_{l})\odot w+R_{q}-R_{l}\big)$
$\quad+\mathord{{\rm div}}\big(w\odot(z_{q+1}-z_{l})+u_{q}\odot\overline{z}+\overline{z}\otimes\overline{z}\big),$
[…]
(4.2) $-\nabla p_{q+1}+\mathord{{\rm div}}(R_{q+1})$ $=\partial_{t}w+u_{l}\cdot\nabla w+\mathord{{\rm div}}\Big(w_{o}\otimes w_{o}-c_{0}\Big)+w\cdot\nabla u_{l}$
$\quad+\mathord{{\rm div}}\Big(w_{c}\otimes w+w_{o}\otimes w_{c}+w\odot(v_{q}-v_{l})+R_{q}-R_{l}\Big)$
[…]
$\quad+\nabla\Big(\frac{2}{3}w\cdot(z_{q+1}-z_{l})+\frac{2}{3}u_{q}\cdot\overline{z}+\frac{|\overline{z}|^{2}}{3}\Big)-\nabla p_{q}.$
Notice that the last line in the expression above is a gradient; hence, we define:
(4.3) $\overline{p}:=p_{q+1}-p_{q}:=-\Big(\frac{2}{3}w\cdot(z_{q+1}-z_{l})+\frac{2}{3}u_{q}\cdot\overline{z}+\frac{|\overline{z}|^{2}}{3}\Big).$
Therefore, in order to enforce (4.2) we can define $R_{q+1}$ as follows:
(4.4) $R_{q+1}$ $:=R_{A}+R_{T}+R_{S}+\frac{2}{3}f(t)\mathrm{Id},$
where $f$ is an arbitrary space-independent scalar function and
(4.5) $R_{A}$ $:=\mathcal{R}\Big[\partial_{t}w+u_{l}\cdot\nabla w+w\cdot\nabla u_{l}+\mathord{{\rm div}}\Big(w_{o}\otimes w_{o}-c_{0}\Big)\Big];$
(4.6) $R_{T}$ $:=w_{c}\otimes w+w_{o}\otimes w_{c}+w\odot(v_{q}-v_{l})+R_{q}-R_{l};$
[…]
In (4.5) above, the operator $\mathcal{R}$ denotes the anti-divergence operator (cf. for instance [DLS13, Definition 4.2]), which to a given smooth periodic vector field $v$ associates a field $\mathcal{R}v$ taking values in the space of symmetric and traceless matrices and satisfying $\mathord{{\rm div}}\mathcal{R}v=v-\fint_{\mathbb{T}^{3}}v$. In fact, the antidivergence stress $R_{A}$ denotes the part of the Reynolds stress $R_{q+1}$ containing all the terms that we will estimate through properties of the the antidivergence operator $\mathcal{R}$. Similarly, the trace stress $R_{T}$ contains all the contributions of terms having trace (except for the term $\frac{2}{3}f(t)\mathrm{Id}$), and the stochastic stress $R_{S}$ contains all the terms with an explicit dependence of the stochastic process $z$. Our particular $f$ entering in the convex integration scheme will be defined by (5.20) below. It will depend on the flux current $\phi$ and it will serve the purpose of canceling out the space average of several terms appearing in the equation (5.3) for the new current $\phi_{q+1}$. In passing, we record the following expression for the trace of $R_{q+1}$, that we shall use later:
(4.8) $\operatorname{Tr}\big(R_{q+1}\big)=\operatorname{Tr}\big(R_{q}-R_{l}\big)+|w_{c}|^{2}+2w_{c}\cdot w_{o}+2(v_{q}-v_{l})\cdot w+2f(t).$
4.1. Estimates on the Reynolds stress
We denote $D_{t,q+1}:=\partial_{t}+u_{q+1}\cdot\nabla$. Our main proposition on the Reynolds stress is the following: morally speaking, taking either a space derivative or and advective derivative of the new Reynolds stress $R_{q+1}$ “costs” a factor $\lambda_{q+1}$.
Proposition 4.1.
There exists $b_{0}:=b_{0}(\alpha)>1$ such that, for any $b\in(1,b_{0})$, there are parameters $\gamma:=(b-1)^{2}$, $a_{0}>1$ and $n_{0},h_{0}\in\mathbb{N}$ sufficiently large with the following property. For every $a>a_{0}$, $t\leq\mathfrak{t}$, and $N\leq n_{0}+3$ the following estimates hold:
(4.9) $\left\|\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\right)w(t)\right\|_{N}$ $\leq\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2},$
(4.10) $\left\|D_{t,q+1}\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\right)w(t)\right\|_{N-1}$ $\leq\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2},$
(4.11) $\left\|R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\right\|_{N}$ $\leq\lambda_{q+1}^{N-2\gamma}\delta_{q+2},$
(4.12) $\left\|D_{t,q+1}\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\,\right)\right\|_{N-1}$ $\leq\lambda_{q+1}^{N-2\gamma}\delta_{q+2}\delta_{q+1}^{1/2}.$
Before proving the previous proposition, let us introduce the following parameters:
(4.13) $n_{0}$ $:=\left\lceil\frac{24b}{b-1}\right\rceil-1,\quad h_{0}:=\left\lceil\frac{8b}{b-1}\right\rceil.$
They will play a crucial role when perfoming estimates involving quantities in which we need to apply the antidivergence operator $\mathcal{R}$, more specifically in the bounds (A.19), (A.46), (A.60) from the Appendix. The values of $n_{0},h_{0}$ ensure that
(4.14) $\frac{\rho}{\mu}\lambda_{q}\delta_{q}^{1/2}(\mu\lambda_{q+1})^{-n_{0}-1}$ $\lesssim\lambda_{q+1}^{-2\gamma}\delta_{q+2}^{3/2},$
(4.15) $\frac{\rho}{\epsilon}\lambda_{q}\delta_{q}^{1/2}(l\lambda_{q+1})^{-h_{0}}$ $\lesssim\lambda_{q+1}^{-1-2\gamma}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2}.$
4.1.1. Proof of Proposition˜4.1: bounds on $R_{T},R_{A},R_{S}$.
As a first step towards proving Proposition˜4.1, we separately give bounds on the stress tensors $R_{T},R_{A},R_{S}$ appearing in the decomposition of the new Reynolds stress $R_{q+1}$.
Lemma 4.2.
Let $R_{T}$ be defined by (4.6). For $N\leq n_{0}+3$, it holds for any $t\leq\mathfrak{t}$ that
(4.16) $\|R_{T}(t)\|_{N}$ $\lesssim l^{1-N}\lambda_{q}^{1-\gamma}\delta_{q+1}(l^{-1}\delta_{q}^{1/2}l_{\rm temp}+1)$
(4.17) $\quad+\lambda_{q+1}^{N}\rho^{1/2}\left(\frac{\rho^{1/2}\mu^{-1}}{\lambda_{q+1}}+l\lambda_{q}\delta_{q}^{1/2}\right);$
(4.18) $\|D_{t,q+1}R_{T}(t)\|_{N-2}$ $\lesssim\lambda_{q}^{N-1-\gamma}\delta_{q+1}\delta_{q}^{1/2}+\lambda_{q+1}^{N-2}\lambda_{q}^{1-\gamma}\delta_{q+1}\rho^{1/2}+l^{1-N}\lambda_{q}^{-\gamma}\delta_{q+1}l_{\rm temp}^{-1}$
(4.19) $\quad+\lambda_{q+1}^{N-2}\rho^{1/2}\left[\lambda_{q+1}\rho^{1/2}\left(\frac{\rho^{1/2}\mu^{-1}}{\lambda_{q+1}}+l\lambda_{q}\delta_{q}^{1/2}\right)\right].$
Proof.
We start by estimating $R_{q}-R_{l}=(R_{q}-\mathbf{P}_{\leq l^{-1}}R_{q})+(\mathbf{P}_{\leq l^{-1}}R_{q}-R_{l}).$ Notice that (3.36) implies that
(4.20) $\mathbf{P}_{\leq l^{-1}}R_{q}-R_{l}(t,x)=\int_{-l_{\rm temp}}^{0}ds\,\eta_{l_{\rm temp}}(s)\int_{s}^{0}(D_{t,l}\mathbf{P}_{\leq l^{-1}}R_{q})(t+r,\mu(t;t+r,x))dr.$
Moreover, due to (2.14), (2.16), and [DLK23, Lemma A.4], we have the following inequality:
(4.21) $\|D_{t,l}\mathbf{P}_{\leq l^{-1}}R_{q}\|_{N}$ $\leq\|\mathbf{P}_{\leq l^{-1}}(D_{t,q}R_{q})\|_{N}+\|[u_{l}\cdot\nabla,\mathbf{P}_{\leq l^{-1}}]R_{q}\|_{N}+\|\mathbf{P}_{\leq l^{-1}}((u_{q}-u_{l})\cdot\nabla R_{q})\|_{N}$
(4.22) $\lesssim l^{-N}\lambda_{q}^{1-\gamma}\delta_{q+1}\delta_{q}^{1/2}+l^{1-N}\|u_{q}\|_{1}\lambda^{1-\gamma}_{q}\delta_{q+1}\lesssim l^{-N}\lambda_{q}^{1-\gamma}\delta_{q+1}\delta_{q}^{1/2}.$
Therefore, by (4.20), (4.21), [BD15+, Proposition C.1], and (A.3) we obtain the bound
(4.23) $\|\mathbf{P}_{\leq l^{-1}}R_{q}(t)-R_{l}(t)\|_{N}\lesssim l^{-N}\lambda_{q}^{1-\gamma}\delta_{q+1}\delta_{q}^{1/2}l_{\rm temp}.$
On the other hand, Bernstein inequality and $\lambda_{q}<l^{-1}$ imply that
[…]
$\lesssim l^{1-N}\lambda_{q}^{1-\gamma}\delta_{q+1}.$
As for the material derivative
(4.25) $\|D_{t,q+1}(R_{q}-R_{l})(t)\|_{N-2}$ $\leq\|D_{t,q+1}R_{q}(t)\|_{N-2}+\|D_{t,q+1}R_{l}(t)\|_{N-2}$
$\leq\|D_{t,q}R_{q}(t)\|_{N-2}+\|(u_{q+1}-u_{q})\cdot\nabla R_{q}(t)\|_{N-2}$
$+\|D_{t,l}R_{l}(t)\|_{N-2}+\|(u_{q+1}-u_{l})\cdot\nabla R_{l}(t)\|_{N-2}$
$\lesssim\lambda_{q}^{N-1-\gamma}\delta_{q+1}\delta_{q}^{1/2}+\lambda_{q+1}^{N-2}\lambda_{q}^{1-\gamma}\delta_{q+1}\rho^{1/2}+l^{2-N}\lambda_{q}^{-\gamma}\delta_{q+1}l_{\rm temp}^{-1},$
where the last inequality is due to (2.16), (3.50), (3.51), the bounds $l<l_{\rm temp}$ and $l_{\rm temp}\|u_{q}\|_{1}\leq 1$, as well as the inequality $\|R_{l}(t)\|_{N}\lesssim l^{-N}\lambda_{q}^{-\gamma}\delta_{q+1}$ valid for $N\geq 0$.
Now we estimate $R_{T}^{\prime}:=R_{T}-(R_{q}-R_{l})=w_{c}\otimes w+w_{o}\otimes w_{c}+w\odot(v_{q}-v_{l})$. The bounds (3.47), (3.48), and (2.14) imply
(4.26) $\|R_{T}^{\prime}(t)\|_{N}\lesssim\lambda_{q+1}^{N}\rho^{1/2}\left(\frac{\rho^{1/2}\mu^{-1}}{\lambda_{q+1}}+l\lambda_{q}\delta_{q}^{1/2}\right).$
Instead, the material derivative is controlled thanks to bounds (3.47), (3.48), (3.61), (3.51), as well as $l\lambda_{q}\delta_{q}^{1/2}\leq\rho^{1/2}$ and $\epsilon^{-1}\lesssim\lambda_{q+1}\rho^{1/2}$, giving
(4.27) $\|D_{t,q+1}R_{T}^{\prime}\|_{N-2}$ $\leq\|D_{t,l}R_{T}^{\prime}\|_{N-1}+\|(u_{q+1}-u_{l})\cdot\nabla R_{T}^{\prime}\|_{N-2}$
$\lesssim\lambda_{q+1}^{N-2}\rho^{1/2}\left[\lambda_{q+1}\rho^{1/2}\left(\frac{\rho^{1/2}\mu^{-1}}{\lambda_{q+1}}+l\lambda_{q}\delta_{q}^{1/2}\right)\right].$
Lemma 4.3.
Let $R_{A}$ be defined by (4.5). For $N\leq n_{0}+3$ and $t\leq\mathfrak{t}$ we have
(4.28) $\|R_{A}(t)\|_{N}$ $\lesssim\lambda_{q+1}^{N}\left[\lambda_{q+1}^{-1}+(\lambda_{q+1}\mu)^{-n_{0}-1}\right]\frac{\rho}{\mu},$
(4.29) $\|D_{t,q+1}R_{A}(t)\|_{N-2}$ $\lesssim\lambda_{q+1}^{N-2}\left[(\epsilon\lambda_{q+1})^{-1}+\lambda_{q+1}^{-n_{0}}\mu^{-n_{0}-1}+(l\lambda_{q+1})^{-h_{0}}\lambda_{q+1}^{2}\lambda_{q}\delta_{q}^{1/2}\right]\frac{\rho}{\mu}$
Proof.
Let us preliminarily observe that
(4.30) $R_{A}$ $=\mathcal{R}\left(\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\sum_{m\in\mathbb{Z}}a_{m,k}e^{i\lambda_{q+1}\xi_{m}\cdot k}\right),$
(4.31) $a_{m,k}:$ $=\mathord{{\rm div}}(c_{m,k})+D_{t,l}(b_{m,k}+\lambda_{q+1}^{-1}e_{m,k})+(b_{m,k}+\lambda_{q+1}^{-1}e_{m,k})\cdot\nabla u_{l}.$
[…]
(4.32) $\|D_{t,l}\mathord{{\rm div}}(c_{m,k})\|_{j}$ $\leq\|D_{t,l}c_{m,k}\|_{j+1}+\|\nabla u_{l}:\nabla c_{m,k}^{T}\|_{j}\lesssim\mu^{-j}\frac{\rho}{\mu\,\epsilon},$
(4.33) $\left\|D_{t,l}\left[(b_{m,k}+\lambda_{q+1}^{-1}e_{m,k})\cdot\nabla u_{l}\right]\right\|_{j}$ $\lesssim\mu^{-j}\rho^{1/2}(\epsilon^{-1}\lambda_{q}\delta_{q}^{1/2}+\lambda_{q}^{2}\delta_{q}).$
Lemma 4.4.
Let $R_{S}$ be defined by (4.7). For $N\leq n_{0}+3$ and $t\leq\mathfrak{t}$ we have
(4.34) $\|R_{S}(t)\|_{N}$ $\lesssim i_{q}^{1/2-\delta}\begin{cases}1,\,\quad&N=0\\ \lambda_{q+1}^{N}\delta_{q+1}^{1/2},\quad&N>0\end{cases},$
(4.35) $\|D_{t,q+1}R_{S}(t)\|_{N-2}$ $\lesssim i_{q+1}^{-1/2-\delta}\begin{cases}1\vee\delta_{q+1}^{1/2},\quad&N=2\\ \lambda_{q+1}^{N-2}\delta_{q+1}^{1/2},\quad&N>2\end{cases}$
(4.36) $\quad+i_{q}^{1/2-\delta}\left[\lambda_{q+1}\rho(\delta_{q+1}\vee\delta_{q+1}^{1/2})\right]\lambda_{q+1}^{N-2}.$
Proof.
We recall that, by definition:
[…]
Therefore, the desired bounds follow from Lemmas˜3.8, 3.3 and 3.5, as well as
(4.38) $\|D_{t,q+1}u_{q}\|_{N}\leq\|D_{t,q}u_{q}\|_{N}+\|(u_{q+1}-u_{q})\cdot\nabla u_{q}\|_{N}\lesssim i_{q}^{-1/2-\delta}+\lambda_{q+1}^{N}\lambda_{q}\,\delta_{q}.$
4.1.2. Proof of Proposition˜4.1: Conclusion
We are ready to show Proposition˜4.1. First, bounds (4.11) and (4.12) follow directly from previous Lemmas˜4.2, 4.3 and 4.4, after some some standard computations. Instead, the bounds (4.9) and (4.10) are due to Lemmas˜4.2, 4.3 and 4.4 and Proposition˜3.3 combined with
(4.39) $\|Xw(t)\|_{N}\lesssim\sum_{j=0}^{N}\|X(t)\|_{j}\|w(t)\|_{N-j}$
applied with either $X(t):=R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}$ or $X(t):=D_{t,q+1}\left(R_{q+1}(t)-\frac{2}{3}f(t)\mathrm{Id}\right).$
4.2. Estimates on the pressure
Our next proposition concerns the new pressure $p_{q+1}$. We have the following:
Proposition 4.5.
There exists $b_{0}:=b_{0}(\alpha)>1$ such that, for any $b\in(1,b_{0})$, there is $a_{0}>1$ such that for every $a>a_{0}$, $t\leq\mathfrak{t}$ the following estimates hold:
(4.40) $\|p_{q+1}(t)\|_{N}$ $\leq\overline{M}\lambda_{q+1}^{N}\delta_{q+1},\quad\|D_{t,q+1}p_{q+1}(t)\|_{N-2}\leq\overline{M}\lambda_{q+1}^{N-1}\delta_{q+1}^{3/2},$ $\text{ for }N\in\{1,2,\dots,n_{0}+3\},$
(4.41) $\|p_{q+1}(t)-p_{q}(t)\|_{N}$ $\leq\overline{M}\lambda_{q+1}^{N}\delta_{q+1}$ $\text{for }N\in\{0,1\}.$
Proof.
Since $p_{q+1}-p_{q}=-\Big(\frac{2}{3}w\cdot(z_{q+1}-z_{l})+\frac{2}{3}u_{q}\cdot\overline{z}+\frac{|\overline{z}|^{2}}{3}\Big)$ by construction, the estimates on $\|p_{q+1}\|_{N}$ and $\|p_{q+1}-p_{q}\|_{N}$ can be easily obtained by mimicking those on $R_{S}$. Analogously,
(4.42) $\|D_{t,q+1}p_{q+1}\|_{N-2}$ $\lesssim\lambda_{q+1}^{N-1}\delta_{q+2}\delta_{q+1}^{1/2}+\|D_{t,q}p_{q}\|_{N-2}+\|(u_{q+1}-u_{q})\cdot\nabla p_{q}\|_{N-2}$
(4.43) $\lesssim\lambda_{q+1}^{N-1}\delta_{q+2}\delta_{q+1}^{1/2}+\lambda_{q}^{N-1}\delta_{q}^{3/2}+\lambda_{q+1}^{N-2}\lambda_{q}\delta_{q}\delta_{q+1}^{1/2}\leq\overline{M}\lambda_{q+1}^{N-1}\delta_{q+1}^{3/2},$
where the last inequality is due to picking $b_{0}$ and $a_{0}$ properly. ∎
5. Construction of the new current
In the previous sections we have constructed the new velocity $v_{q+1}$, pressure $p_{q+1}$, and Reynolds stress $R_{q+1}$. Moreover, the noise $z_{q+1}$ is given by (2.11). The new current $\phi_{q+1}$ is defined by imposing that $(v_{q+1},p_{q+1},R_{q+1},\phi_{q+1},z_{q+1})$ satisfies the relaxed local energy equality in (2.13), namely
(5.1) $\partial_{t}$ $\frac{|v_{q+1}|^{2}}{2}+v_{q+1}\cdot\nabla p_{q+1}+v_{q+1}\cdot\mathord{{\rm div}}\big((v_{q+1}+z_{q+1})\otimes(v_{q+1}+z_{q+1})\big)=-E^{\prime}$
(5.2) $+\frac{1}{2}D_{t,q+1}\big(\operatorname{Tr}(R_{q+1})\big)+\mathord{{\rm div}}\big(R_{q+1}v_{q+1}\big)+R_{q+1}:\nabla z_{q+1}^{T}+\mathord{{\rm div}}(\phi_{q+1}).$
Since $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$ solves itself the relaxed local energy inequality, we have
(5.3) $\partial_{t}$ $\frac{|v_{q+1}|^{2}}{2}+v_{q+1}\cdot\nabla p_{q+1}+v_{q+1}\cdot\mathord{{\rm div}}\big(u_{q+1}\otimes u_{q+1}\big)=-E^{\prime}$
(5.4) $+\mathord{{\rm div}}(R_{q+1}v_{q+1})+R_{q+1}:\nabla z_{q+1}^{T}$
(5.5) $+D_{t,q}\left(\frac{\operatorname{Tr}(R_{q})+|w|^{2}+2(v_{q}-v_{l})\cdot w}{2}\right)+\mathord{{\rm div}}\left(\phi_{q}+\frac{|w|^{2}}{2}w-R_{q+1}w+(\overline{z}\otimes v_{q})w\right)$
(5.6) $+\mathord{{\rm div}}\left((R_{q}+w\otimes w-R_{q+1})(v_{q}-v_{l})+\left[\overline{z}\otimes\overline{z}+u_{q}\odot\overline{z}\right]v_{l}\right)$
(5.7) $+\mathord{{\rm div}}\left(\left(\frac{|v_{q}|^{2}+|w|^{2}}{2}\right)\overline{z}+\frac{|v_{q}-v_{l}|^{2}}{2}w\right)$
(5.8) $+\left[(w_{o}\otimes w_{o}-c_{0})-\mathcal{R}(\mathord{{\rm div}}(w_{o}\otimes w_{o}-c_{0})+D_{t,l}w+w\cdot\nabla u_{l})\right]:\nabla(v_{l}+z_{q})^{T}-R_{q+1}:\nabla\overline{z}^{T}$
[…]
(5.10) $-(z_{q+1}-z_{l})\cdot\mathord{{\rm div}}(w\otimes v_{l})+(v_{q}-v_{l})\cdot\mathord{{\rm div}}(w\otimes\overline{z})+w\cdot\mathord{{\rm div}}\Big\{(p_{q+1}-p_{l})\mathrm{Id}+R_{l}$
(5.11) $+R_{comm}-z\otimes v_{l}+(z_{q+1}-z_{l})\otimes z_{q}+z_{l}\otimes(z_{q}-z_{l})-v_{l}\otimes(z_{l}-z_{q})+u_{q+1}\otimes\overline{z}\Big\},$
where $R_{comm}:=u_{l}\otimes u_{l}-\mathbf{P}_{\leq l^{-1}}(u_{q}\otimes u_{q})$, and we have used multiple times the divergence-free property of all the vector fields at stake, as well as the following standard algebraic identities valid for generic matrices $M$ and vector fields $v,u$:
(5.12) $\mathord{{\rm div}}(Mv)$ $=\mathord{{\rm div}}(M)\cdot v+M:\nabla v^{T},$
(5.13) $u\otimes v:\nabla w^{T}$ $=v\cdot\mathord{{\rm div}}(u\otimes w).$
Equation (4.8) implies that
(5.14) $D_{t,q}\Big(\frac{\operatorname{Tr}(R_{q})+|w|^{2}+2(v_{q}-v_{l})\cdot w}{2}\Big)$ $=D_{t,q}\left(\frac{\operatorname{Tr}(R_{q+1})+\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})}{2}\right)-\partial_{t}f$
(5.15) $=D_{t,q+1}\left(\frac{\operatorname{Tr}(R_{q+1})}{2}\right)-\mathord{{\rm div}}\left((w+\overline{z})\frac{\operatorname{Tr}(R_{q+1})}{2}\right)$
(5.16) $\quad+\mathord{{\rm div}}\left(\mathcal{R}\left(D_{t,l}\frac{\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})}{2}\right)\right)$
(5.17) $\quad+\fint_{\mathbb{T}^{3}}D_{t,l}\frac{\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})}{2}\,dx$
(5.18) $\quad+\mathord{{\rm div}}\left((v_{q}-v_{l})\frac{\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})}{2}\right)-\partial_{t}f.$
Therefore, (5.1) holds for $\phi_{q+1}$ and $f(t)$ given by:
(5.19) $\phi_{q+1}$ $:=\phi_{A}+\phi_{D_{1}}+\phi_{D_{2}}+\phi_{S_{1}}+\phi_{S_{2}},$
(5.20) $f_{q+1}(t)$ $:=f_{A}(t)+f_{D_{2}}(t)+f_{S_{2}}(t):=\int_{0}^{t}\fint_{\mathbb{T}^{3}}\left(\varphi_{A}(x,s)+\varphi_{D_{2}}(x,s)+\varphi_{S_{2}}(x,s)\right)\,dx\,ds,$
where the terms at the right-hand sides above are defined as follows: the antidivergence current
(5.21) $\phi_{A}$ $:=\overline{\mathcal{R}}\varphi_{A},$
(5.22) $\varphi_{A}$ $:=D_{t,l}\frac{\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})}{2}+\mathord{{\rm div}}\left(\phi_{l}+\frac{|w_{o}|^{2}}{2}w_{o}\right)+\left(w_{o}\otimes w_{o}-c_{0}\right):\nabla(v_{l}+z_{q})^{T}$
[…]
(5.24) $\quad+w\cdot\mathord{{\rm div}}\left(\mathbf{P}_{\leq l^{-1}}R_{q}+R_{comm}-\overline{z}\otimes v_{l}+(z_{q+1}-z_{l})\otimes z_{q}\right)$
(5.25) $\quad+w\cdot\mathord{{\rm div}}(u_{l}\otimes(z_{q}-z_{l})+u_{q+1}\otimes\overline{z})+w\cdot\nabla(p_{q}-p_{l}),$
contains all the terms that will be treated by means of the properties of the antidivergence operator $\overline{\mathcal{R}}:=-\nabla(-\Delta)^{-1}$, that to smooth scalar fields $g$ associates a vector field satisfying $\mathord{{\rm div}}\overline{\mathcal{R}}g=g-\fint_{\mathbb{T}^{3}}g$; the stochastic currents
(5.26) $\phi_{S_{1}}$ $:=\left(\frac{\operatorname{Tr}(R_{q+1}-\frac{3}{2}f\,\mathrm{Id})}{2}+\frac{|v_{q}|^{2}+|w|^{2}}{2}\right)\overline{z}+(\overline{z}\otimes v_{q})w+(\overline{z}\otimes\overline{z}+u_{q}\odot\overline{z})v_{l}+\overline{p}\,w\,$,
(5.27) $\phi_{S_{2}}$ $:=\overline{\mathcal{R}}\varphi_{S_{2}},$
(5.28) $\varphi_{S_{2}}$ $:=R_{A}:\nabla(z_{q}-z_{l})^{T}+\left(u_{q}\,\mathring{\odot}\,\overline{z}+\overline{z}\otimes\overline{z}\right):\nabla z_{q}^{T}$
(5.29) $\quad+v_{q}\cdot\mathord{{\rm div}}(u_{q}\otimes\overline{z}+\overline{z}\otimes z_{q+1})-\left(R_{q+1}-\frac{3}{2}f\,\mathrm{Id}\right):\nabla\overline{z}^{T},$
containing all the terms with explicit dependence of the noise $z$; and the deterministic currents
(5.30) $\phi_{D_{1}}$ $:=w\frac{\operatorname{Tr}(R_{q+1}-\frac{3}{2}f\,\mathrm{Id})}{2}+(v_{q}-v_{l})\frac{\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})}{2}+\phi_{q}-\phi_{l}$
(5.31) $+\frac{|w|^{2}}{2}w-\frac{|w_{o}|^{2}}{2}w_{o}-\left(R_{q+1}-\frac{3}{2}f\,\mathrm{Id}\right)w+\frac{|v_{q}-v_{l}|^{2}}{2}w$
(5.32) $+\left(R_{q}-w\otimes w-R_{q+1}-\rho\mathrm{Id}+\frac{3}{2}f\,\mathrm{Id}\right)(v_{q}-v_{l})\,$,
(5.33) $\phi_{D_{2}}$ $:=\overline{\mathcal{R}}\varphi_{D_{2}},\quad\varphi_{D_{2}}:=R_{A}:\nabla u_{l}^{T},$
containing all the remaining terms that would be there even without external random forcing. Obviously, if $z\neq 0$ then the deterministic currents $\phi_{D_{1}}$ and $\phi_{D_{2}}$ are random objects nonetheless, being so all the other objects involved in the construction. Furthermore, the terms $\phi_{D_{1}}$ and $\phi_{S_{1}}$ can be written in the form $\overline{\mathcal{R}}\varphi_{D_{1}}$ and $\overline{\mathcal{R}}\varphi_{S_{1}}$ respectively, where $\varphi_{D_{1}}=\mathord{{\rm div}}\phi_{D_{1}}$ and $\varphi_{S_{1}}=\mathord{{\rm div}}\phi_{S_{2}}$ have zero space average. On the other hand, the currents $\varphi_{A},\varphi_{S_{2}}$, and $\varphi_{D_{2}}$ have non-zero space average.
5.1. Estimates on the current
Our main proposition on the current is the following Proposition˜5.1. The same caveat as before Proposition˜3.3 applies also here: the proof of this results could be skipped at a first reading. Roughly speaking, taking either a space derivative or an advective derivative of the new current $\phi_{q+1}$ “costs” a factor $\lambda_{q+1}$. On the other hand, the scalar function $f$ and its time derivative go to zero as $\lambda_{q+1}^{-\gamma}\delta_{q+2}$ and $\lambda_{q+1}^{-\gamma}\delta_{q+2}\delta_{q+1}^{1/2}$, respectively.
Proposition 5.1.
There exists $b_{0}:=b_{0}(\alpha)>1$ such that, for any $b\in(1,b_{0})$, there are parameters $\gamma:=(b-1)^{2}$, $a_{0}>1$ and $n_{0},h_{0}\in\mathbb{N}$ sufficiently large with the following property. For every $a>a_{0}$, $t\leq\mathfrak{t}$ and $N\leq 2$ it holds
(5.34) $\|\phi_{q+1}(t)\|_{N}$ $\leq\lambda_{q+1}^{N-3\gamma/2}\delta_{q+2}^{3/2},$
(5.35) $\|D_{t,q+1}\phi_{q+1}(t)\|_{N-1}$ $\leq\lambda_{q+1}^{N-3\gamma/2}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2},$
(5.36) $|f(t)|$ $\leq\frac{1}{100}\lambda_{q+1}^{-\gamma}\delta_{q+2},$
(5.37) $|\partial_{t}f(t)|$ $\leq\frac{1}{100}\lambda_{q+1}^{-\gamma}\delta_{q+2}\delta_{q+1}^{1/2}.$
5.1.1. Proof of Proposition˜5.1: bounds on $\phi_{A}$
Lemma 5.2.
Let $\phi_{A}$, $f_{A}$ be defined by (5.21) and (5.20). For $N\leq 2$ and $t\leq\mathfrak{t}$ we have
(5.38) $\|\phi_{A}(t)\|_{N}$ $\lesssim\lambda_{q+1}^{N}\left[\lambda_{q+1}^{-1}+(\lambda_{q+1}\mu)^{-n_{0}-1}\right]\frac{\delta_{q+1}}{\epsilon},$
(5.39) $\|D_{t,q+1}\phi_{A}(t)\|_{N-1}$ $\lesssim\lambda_{q+1}^{N-1}\left[(\epsilon\lambda_{q+1})^{-1}+\lambda_{q+1}^{-n_{0}}\mu^{-n_{0}-1}+(l\lambda_{q+1})^{-h_{0}}\lambda_{q+1}^{2}\lambda_{q}\delta_{q}^{1/2}\right]\frac{\delta_{q+1}}{\epsilon},$
(5.40) $|f_{A}(t)|+|f^{\prime}_{A}(t)|$ $\lesssim\frac{\delta_{q+1}}{\epsilon}(\lambda_{q+1}\mu)^{-n_{0}-1}.$
Proof.
The result is due to Propositions˜A.3 and A.4 applied to $F=\phi_{A}$. Indeed, recall the decomposition of $\phi_{A}$ given by (5.21). Following the notation of Proposition˜A.3, let us define $a_{m,k}:=a^{1}_{m,k}+a^{2}_{m,k}+a^{3}_{m,k}$ where
(5.41) $a_{m,k}^{1}$ $:=D_{t,l}\left(\frac{\operatorname{Tr}(c_{m,k})}{2}\right)+\mathord{{\rm div}}(d_{m,k})+c_{m,k}:\nabla(v_{l}+z_{q})^{T}+\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\,\mathring{\odot}\,(z_{q+1}-z_{l}):\nabla z_{q}^{T}$
(5.42) $\quad+\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\cdot\mathord{{\rm div}}\left((z_{q+1}-z_{l})\otimes z_{q}\right)-(z_{q+1}-z_{l})\cdot\left[\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\cdot\nabla v_{l}\right]$
(5.43) $\quad+\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\cdot\left[-(\overline{z}\cdot\nabla)v_{l}+u_{l}\cdot\nabla(z_{q}-z_{l})+u_{q+1}\cdot\nabla\overline{z}\right]$
(5.44) $\quad+(v_{q}-v_{l})\cdot\left[\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\cdot\nabla\overline{z}\right],$
(5.45) $a_{m,k}^{2}$ $:=\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\mathord{{\rm div}}(\mathbf{P}_{\leq l^{-1}}R_{q}+R_{comm}),$
(5.46) $a_{m,k}^{3}$ $:=\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\cdot\nabla(p_{q}-p_{l}).$
We intend to apply Proposition˜A.3 with $\lambda=\lambda_{q+1}$, a sequence
(5.47) $s_{j}=\frac{\delta_{q+1}}{\epsilon}\mu^{-j},$
[…]
Let us move to the estimate on $\|D_{t,q+1}\phi_{A}(t)\|_{N-1}$. We apply Proposition˜A.4 with $\lambda,s_{j}$ as above and $c=\epsilon$. In this case, we check the assumptions on $D_{t,l}a_{m,k}^{1}$ by the inductive estimates, Lemma˜3.8, the bound (3.64), Lemma˜3.4, and ˜3.5. In a similar fashion, in order to control the assumptions on $D_{t,l}a_{m,k}^{2}$ we use Bernstein inequality, (3.61) and [DLK23, Lemma A.4] to control, on one hand:
(5.48) $\|D_{t,l}(\mathord{{\rm div}}R_{comm})\|_{N}$ $\lesssim l^{-N}\|D_{t,l}(\mathord{{\rm div}}R_{comm})\|_{0}$
(5.49) $\lesssim\|D_{t,l}(u_{q}-u_{l})\|_{0}\|u_{q}\|_{1}+l\|u_{q}\|_{1}\left[\|D_{t,l}u_{l}\|_{1}+\|u_{q}\|_{1}^{2}\right]+l^{2}\|u_{q}\|_{1}^{2}\|u_{q}\|_{2}$
(5.50) $\lesssim l^{1-N}\left[\lambda_{q}^{3}\delta_{q}^{3/2}+\lambda_{q}\delta_{q}^{1/2}i_{q}^{-1/2-\delta}\right],$
where the second inequality follows by the same computations shown in [DLK23, Lemma 5.2], and on the other
(5.51) $\|D_{t,l}(\mathord{{\rm div}}\mathbf{P}_{\leq l^{-1}}R_{q})\|_{N}$ $\leq\|\mathbf{P}_{\leq l^{-1}}D_{t,l}R_{q}\|_{N+1}+\|[u_{l}\cdot\nabla,\mathbf{P}_{\leq l^{-1}}]R_{q}\|_{N+1}+\|(\nabla u_{l})\nabla(\mathbf{P}_{\leq l^{-1}}R_{q})\|_{N}$
(5.52) $\lesssim l^{-N}\|D_{t,q}R_{q}\|_{1}+\|(u_{q}-u_{l})\cdot\nabla R_{q}\|_{1}+l^{-N}\|u_{q}\|_{1}\|R_{q}\|_{1}$
(5.53) $\lesssim l^{-N}\lambda^{2}_{q}\delta_{q+1}\delta_{q}^{1/2}.$
Finally, as for the terms $D_{t,l}a_{m,k}^{3}$, we have by Lemma˜3.8
(5.54) $\|D_{t,l}a_{m,k}^{3}\|_{j}$ $\lesssim\sum_{h=0}^{j}\left\|D_{t,l}\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\right\|_{j-h}\|p_{q}-p_{l}\|_{h+1}$
(5.55) $+\sum_{h=0}^{j}\left\|\left(b_{m,k}+\frac{e_{m,k}}{\lambda_{q+1}}\right)\right\|_{j-h}\left(\|D_{t,l}(p_{q}-p_{l})\|_{h+1}+\|(\nabla u_{l})\nabla(p_{q}-p_{l})\|_{h}\right)$
(5.56) $\lesssim\lambda_{q}\delta_{q}\rho^{1/2}\epsilon^{-1}\mu^{-j},$
where in the last line we have used the identity
(5.57) $D_{t,l}(p_{q}-p_{l})$ $=D_{t,q}p_{q}-(D_{t,q}p_{q})_{l}+(u_{l}-u_{q})\cdot\nabla p_{q}+(u_{q}\cdot\nabla p_{q})_{l}-u_{l}\cdot\nabla u_{l},$
and the following bound, obtained by [DLK23, Lemma A.3] and Bernstein inequality:
(5.58) $\|D_{t,l}(p_{q}-p_{l})\|_{M}\lesssim l^{1-M}\lambda_{q}^{2}\delta_{q}^{3/2}.$
At last, let us consider $|f_{A}(t)|+|f^{\prime}_{A}(t)|$. Due to the spatial and temporal partitions of unity, arguing as in Proposition˜3.3 it is sufficient to estimate $f_{m,k}(t):=\int_{0}^{t}\int_{\mathbb{T}^{3}}a_{m,k}(x,s)e^{i\lambda_{q+1}\xi_{m}(x,s)\cdot k}\,dx\,ds$. Then, [DLK23, Lemma A.2] implies that
(5.59) $|f_{A}(t)|+|f^{\prime}_{A}(t)|\lesssim\lambda_{q+1}^{-n_{0}+1}(s_{n_{0}+1}+s_{0}\,l^{-n_{0}-1})\lesssim\frac{\delta_{q+1}}{\epsilon}(\lambda_{q+1}\mu)^{-n_{0}-1}.$
5.1.2. Proof of Proposition˜5.1: bounds on $\phi_{D}$
Lemma 5.3.
Let $\phi_{D_{1}}$ be defined by (5.30). For $N\leq 2$ and $t\leq\mathfrak{t}$ we have
(5.60) $\|\phi_{D_{1}}(t)\|_{N}$ $\lesssim\lambda_{q+1}^{N}\left[l\,\lambda_{q}\delta_{q}^{1/2}\rho+(\lambda_{q+1}\mu)^{-1}\rho^{3/2}\right]$
(5.61) $\quad+l^{1-N}\lambda_{q}^{1-\gamma}\delta_{q+1}^{3/2}(l^{-1}\delta_{q}^{1/2}l_{\rm temp}+1)+\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2},$
(5.62) $\|D_{t,q+1}\phi_{D_{1}}(t)\|_{N-1}$ $\lesssim\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2}+l^{1-N}\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}l_{\rm temp}^{-1}$
(5.63) $\quad+\lambda_{q+1}^{N}[l\delta_{q+1}^{3/2}\lambda_{q}\delta_{q}^{1/2}+\lambda_{q+1}^{-1}\lambda_{q}^{1-\gamma}\delta_{q+1}^{3/2}\rho^{1/2}+(\lambda_{q+1}\mu)^{-1}\rho^{2}]$
Proof.
We bound each term in the decomposition (5.30) separately, starting from
(5.64) $\bar{\phi}_{D_{1}}:=w\frac{\operatorname{Tr}(R_{q+1}-\frac{3}{2}f\,\mathrm{Id})}{2}-\left(R_{q+1}-\frac{3}{2}f\,\mathrm{Id}\right)w.$
In this case, the bounds (4.9) and (4.10) imply that
(5.65) $\|\bar{\phi}_{D_{1}}\|_{N}\lesssim\lambda_{q+1}^{N-3\gamma/2}\delta_{q+2}^{3/2},\qquad\|D_{t,q+1}\bar{\phi}_{D_{1}}\|_{N-1}\lesssim\lambda_{q+1}^{N-3\gamma/2}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2}.$
Moving to the other terms, we have by (2.14), Lemma˜3.8, (3.90) and the bound $\mu^{-1}<\lambda_{q+1}$ that
(5.66) $\left\|(v_{q}-v_{l})\frac{\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})}{2}\right\|_{N}\lesssim l\lambda_{q}\delta_{q}^{1/2}\rho\lambda_{q+1}^{N},$
while the advective derivative is controlled by
(5.67) $\left\|D_{t,q+1}\left((v_{q}-v_{l})\frac{\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})}{2}\right)\right\|_{N-1}$ $\lesssim\|D_{t,l}(v_{q}-v_{l})\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})\|_{N-1}$
(5.68) $\quad+\|(v_{q}-v_{l})D_{t,l}\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})\|_{N-1}$
(5.69) $\quad+\left\|((u_{q+1}-u_{l})\cdot\nabla)\left((v_{q}-v_{l})\operatorname{Tr}(w_{o}\otimes w_{o}-c_{0})\right)\right\|_{N-1}$
(5.70) $\lesssim\lambda_{q+1}^{N}l\,\rho^{3/2}\lambda_{q}\delta_{q}^{1/2},$
where in the last line we have used (5.66) above, Lemma˜3.8, the bounds (3.61), (3.51), and Proposition˜3.3. By similar computations, we can control also
(5.71) $\left\|\frac{|v_{q}-v_{l}|^{2}}{2}w\right\|_{N}$ $\lesssim\lambda_{q+1}^{N}l^{2}\lambda_{q}^{2}\delta_{q}\,\rho^{1/2},$
(5.72) $\left\|D_{t,q+1}\left(\frac{|v_{q}-v_{l}|^{2}}{2}w\right)\right\|_{N-1}$ $\lesssim\lambda_{q+1}^{N}l^{2}\lambda_{q}^{2}\delta_{q}\,\rho.$
The next term in the decomposition (5.30) we deal with is $\phi_{q}-\phi_{l}$. In this case, we can repeat the computations reported in Lemma˜4.2, obtaining
(5.73) $\|\phi_{q}-\phi_{l}\|_{N}$ $\lesssim l^{1-N}\lambda_{q}^{1-\gamma}\delta_{q+1}^{3/2}(l^{-1}\delta^{1/2}_{q}l_{\rm temp}+1),$
(5.74) $\|D_{t,q+1}(\phi_{q}-\phi_{l})\|_{N-1}$ $\lesssim\lambda_{q}^{N-\gamma}\delta_{q+1}^{3/2}\delta_{q}^{1/2}+\lambda_{q+1}^{N-1}\lambda_{q}^{1-\gamma}\delta_{q+1}^{3/2}\rho^{1/2}+l^{1-N}\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}l_{\rm temp}^{-1}.$
The cubic term in the velocity perturbation is controlled by means of Proposition˜3.3 as follows:
(5.75) $\left\|\frac{|w|^{2}w-|w_{o}|^{2}w_{o}}{2}\right\|_{N}$ $\lesssim\lambda_{q+1}^{N}(\lambda_{q+1}\mu)^{-1}\rho^{3/2},$
(5.76) $\left\|D_{t,q+1}\left(\frac{|w|^{2}w-|w_{o}|^{2}w_{o}}{2}\right)\right\|_{N-1}$ $\lesssim\lambda_{q+1}^{N}(\lambda_{q+1}\mu)^{-1}\rho^{2}.$
Finally, the remaining term $\tilde{\phi}_{D_{1}}$ can be rewritten as $\tilde{\phi}_{D_{1}}=\bar{R}(v_{q}-v_{l})$, where
(5.77) $\bar{R}$ $:=R_{q}-w\otimes w-R_{q+1}-\rho\mathrm{Id}+\frac{3}{2}f\,\mathrm{Id}$
(5.78) $=-\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}c_{m,k}e^{i\lambda_{q+1}k\cdot\xi_{m}}-R_{A}-R_{S}-w\odot(v_{q}-v_{l}),$
and then we can control it by Lemmas˜3.8, 4.3 and 4.4 as follows
(5.79) $\|\tilde{\phi}_{D_{1}}\|_{N}$ $\lesssim\lambda_{q+1}^{N}l\,\lambda_{q}\delta_{q}^{1/2}\rho,$
whereas the advective derivative can be controlled by using the aforementioned lemmas, Proposition˜3.3, and the estimates (3.51), (3.61)
(5.80) $\|D_{t,q+1}\tilde{\phi}_{D_{1}}\|_{N-1}$ $\leq\left\|(v_{q}-v_{l})D_{t,q+1}(R_{A}+R_{S}+w\odot(v_{q}-v_{l}))\right\|_{N-1}$
(5.81) $\quad+\left\|(v_{q}-v_{l})\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\left(D_{t,l}(c_{m,k})e^{i\lambda_{q+1}k\cdot\xi_{m}}+(u_{q+1}-u_{l})\cdot\nabla(c_{m,k}e^{i\lambda_{q+1}k\cdot\xi_{m}})\right)\right\|_{N-1}$
(5.82) $\quad+\left\|\left(D_{t,l}(v_{q}-v_{l})+(u_{q+1}-u_{l})\cdot\nabla(v_{q}-v_{l})\right)\bar{R}\right\|_{N-1}$
(5.83) $\lesssim\lambda_{q+1}^{N-1}\lambda_{q}\delta_{q}^{1/2}\left(\lambda_{q+1}l\,\delta_{q+1}^{3/2}+\rho^{3/2}\right).$
Lemma 5.4.
Let $\phi_{D_{2}}$, $f_{D_{2}}$ be defined by (5.33) and (5.20). For $N\leq 2$, it holds for any $t\leq\mathfrak{t}$ that
(5.84) $\|\phi_{D_{2}}(t)\|_{N}$ $\lesssim\frac{\rho}{\mu}\lambda_{q+1}^{N}\lambda_{q}\delta_{q}^{1/2}(\lambda_{q+1}^{-2}+\mu^{-n_{0}-1}\lambda_{q+1}^{-n_{0}-1}),$
(5.85) $\|D_{t,q+1}\phi_{D_{2}}(t)\|_{N-1}$ $\lesssim\frac{\rho}{\mu}\lambda_{q+1}^{N-1}\Bigg[\rho^{1/2}\lambda_{q}\delta_{q}^{1/2}(\lambda_{q+1}^{-1}+\lambda_{q+1}(\lambda_{q+1}\mu)^{-n_{0}-1})$
(5.86) $\quad+\left(\frac{\lambda_{q}\delta_{q}^{1/2}\epsilon^{-1}}{\lambda_{q+1}^{2}}+\frac{\lambda_{q}^{2}\delta_{q}\,l}{\lambda_{q+1}}\right)+(\lambda_{q+1}\mu)^{-n_{0}-1}\left(\,\lambda_{q+1}\lambda_{q}\delta_{q}^{1/2}\right)\Bigg],$
(5.87) $|f_{D_{2}}(t)|+|f^{\prime}_{D_{2}}(t)|$ $\lesssim\frac{\rho}{\mu}\lambda_{q}\delta_{q}^{1/2}(\lambda_{q+1}\mu)^{-n_{0}-1}.$
Proof.
Based on the decompositions (5.33) and (4.5), we have that
(5.88) $\phi_{D_{2}}=\overline{\mathcal{R}}\left[\mathcal{R}\left(G_{1}+G_{2}+G_{3}\right):\nabla u_{l}^{T}\right],$
where we have denoted $G_{1}:=\mathord{{\rm div}}(w_{o}\otimes w_{o}-c_{0})$, $G_{2}:=D_{t,l}w$, and $G_{3}:=w\cdot\nabla u_{l}$. Notice that, for every $i=1,2,3$, $\mathcal{R}\mathbf{P}_{\geq 3\lambda_{q+1}/8}G_{i}:\nabla u_{l}^{T}$ has frequencies of size at least $3\lambda_{q+1}/8-l^{-1}\gtrsim\lambda_{q+1}$ and $\mathcal{R}\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}:\nabla u_{l}^{T}$ has frequencies of size at most $3\lambda_{q+1}/8+l^{-1}\lesssim\lambda_{q+1}$. Therefore, by the bound on the antidivergence $\|\overline{\mathcal{R}}\mathbf{P}_{\gtrsim\lambda_{q+1}}f\|_{N}\lesssim\lambda_{q+1}^{-1}\|f\|_{N}$ we have
(5.89) $\left\|\overline{\mathcal{R}}\left[\mathcal{R}\left(G_{i}\right):\nabla u_{l}^{T}\right]\right\|_{N}$ $\leq\left\|\overline{\mathcal{R}}\left[\mathcal{R}\mathbf{P}_{\geq 3\lambda_{q+1}/8}G_{i}:\nabla u_{l}^{T}\right]\right\|_{N}+\left\|\overline{\mathcal{R}}\left[\mathcal{R}\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}:\nabla u_{l}^{T}\right]\right\|_{N}$
(5.90) $\lesssim\frac{1}{\lambda_{q+1}}\left\|\mathcal{R}\mathbf{P}_{\geq 3\lambda_{q+1}/8}G_{i}:\nabla u_{l}^{T}\right\|_{N}+{}\lambda_{q+1}^{N}\left\|\mathcal{R}\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}:\nabla u_{l}^{T}\right\|_{0}$
(5.91) $\lesssim\frac{1}{\lambda_{q+1}^{2}}\sum_{j=0}^{N}\|G_{i}\|_{j}\,l^{N-j}\|u_{q}\|_{1}+{\lambda_{q+1}^{N}}\|\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}\|_{0}\|u_{q}\|_{1}.$
Now, recalling the definition on the antidivergence stress $R_{A}$ given by (4.5), we observe that each $G_{i}$ has already been treated in the proof of Lemma˜4.3. Therefore, the estimates therein and (3.90) imply that
(5.92) $\|G_{i}\|_{j}\lesssim\frac{\rho}{\mu}\lambda_{q+1}^{j}.$
Instead, (A.20) implies that $\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}=-\sum_{m,k}\epsilon_{n_{0}}^{\lambda}(k\cdot\xi_{m},a_{m,k}^{i})e^{i\lambda_{q+1}k\cdot\xi_{m}}$, so that by (A.38)
(5.93) $\|\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}\|_{0}\lesssim\frac{\rho}{\mu}\mu^{-n_{0}-1}\lambda_{q+1}^{-n_{0}-1}.$
Therefore, (5.89), (5.92), and (5.93) produce the following bound
(5.94) $\|\phi_{D_{2}}\|_{N}\lesssim\frac{\rho}{\mu}\lambda_{q+1}^{N}\lambda_{q}\delta_{q}^{1/2}(\lambda_{q+1}^{-2}+\mu^{-n_{0}-1}\lambda_{q+1}^{-n_{0}-1}).$
Now, let us turn to Eq.˜5.85. Denote $H_{i}:=\mathcal{R}\mathbf{P}_{\geq 3\lambda/8}G_{i}:\nabla u_{l}^{T},$ then
(5.95) $\|D_{t,l}\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$ $\leq\|\mathcal{R}\left(\mathbf{P}_{\geq\lambda_{q+1}}D_{t,l}H_{i}+[u_{l}\cdot\nabla,\mathbf{P}_{\geq\lambda_{q+1}}]H_{i}\right)\|_{N-1}$
(5.96) $\quad+\|[u_{l}\cdot\nabla,\mathcal{R}]\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$
(5.97) $\lesssim\frac{1}{\lambda_{q+1}}\|D_{t,l}H_{i}+[u_{l}\cdot\nabla,\mathbf{P}_{\geq\lambda_{q+1}}]H_{i}\|_{N-1}+\|[u_{l}\cdot\nabla,\mathcal{R}]\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}$
(5.98) $\lesssim\frac{1}{\lambda_{q+1}}\|D_{t,l}H_{i}\|_{N-1}+\lambda_{q+1}^{N-3}\|u_{q}\|_{1}\|H_{i}\|_{1}+\sum_{j_{1}+j_{2}\leq N-1}l\,\|\nabla u_{l}\|_{j_{1}}\|H_{i}\|_{j_{2}},$
where the first inequality is due to $\mathcal{R}$, the second to a trivial adaptation of [DLK23, Lemma A.4], ˜A.5 and $(l\lambda_{q+1})^{-h_{0}}\lambda^{2}_{q+1}\leq l$. By a similar reasoning as above, we can estimate $\|D_{t,l}(\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}G_{i})\|_{M}$. Indeed, it holds:
(5.99) $\|D_{t,l}(\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}G_{i})\|_{M}\lesssim\frac{1}{\lambda_{q+1}}\|D_{t,l}G_{i}\|_{M}+\lambda_{q+1}^{M-2}\|u_{l}\|_{1}\|G_{i}\|_{1}+\sum_{M_{1}+M_{2}\leq M}l\|\nabla u_{l}\|_{M_{1}}\|G_{i}\|_{M_{2}}.$
Thanks to the previous inequality, Eqs.˜3.64 and 4.32, we have that
(5.100) $\|D_{t,l}H_{i}\|_{N-1}$ $\leq\|D_{t,l}(\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}G_{i}):\nabla u_{l}^{T}\|_{N-1}+\|(\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}G_{i}):D_{t,l}\nabla u_{l}^{T}\|_{N-1}$
(5.101) $\lesssim\lambda^{N-1}_{q+1}\frac{\rho}{\mu}\left(\frac{\lambda_{q}\delta_{q}^{1/2}\epsilon^{-1}}{\lambda_{q+1}^{2}}+\frac{\lambda_{q}^{2}\delta_{q}\,l}{\lambda_{q+1}}\right)$
Instead by standard computations, it holds true
(5.102) $\|H_{i}\|_{M}\lesssim\lambda_{q+1}^{M-1}\frac{\rho}{\mu}\lambda_{q}\delta_{q}^{1/2}.$
Therefore, Eqs.˜5.95, 5.100 and 5.102 imply that
(5.103) $\|D_{t,l}\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}\lesssim\lambda^{N-1}_{q+1}\frac{\rho}{\mu}\left(\frac{\lambda_{q}\delta_{q}^{1/2}\epsilon^{-1}}{\lambda_{q+1}^{2}}+\frac{\lambda_{q}^{2}\delta_{q}\,l}{\lambda_{q+1}}\right).$
Instead, Bernstein inequality, Eqs.˜5.93 and A.60 imply that
(5.104) $\|D_{t,l}\mathcal{R}(\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{N-1}$ $\lesssim\lambda_{q+1}^{N-1}\big[\|D_{t,l}(\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{0}$
(5.105) $+\|[u_{l}\cdot\nabla,\mathcal{R}](\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{0}\big]$
(5.106) $\lesssim\lambda_{q+1}^{N-1}\Big[\big(\|D_{t,l}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}\|_{0}+\|[u_{l}\cdot\nabla,\mathcal{R}](\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i})\|_{0}\big)\|u_{l}\|_{1}$
(5.107) $+\|\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}\|_{0}\|D_{t,l}\nabla u_{l}\|_{0}+\lambda_{q+1}\,\|u_{l}\|_{0}\|u_{l}\|_{1}\|\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}\|_{0}\Big]$
(5.108) $\lesssim\lambda_{q+1}^{N-1}\frac{\rho}{\mu}(\lambda_{q+1}\mu)^{-n_{0}-1}\left[\,\lambda_{q+1}\lambda_{q}\delta_{q}^{1/2}\right]$
At last, Eqs.˜3.51 and 5.84
(5.109) $\|(u_{q+1}-u_{l})\cdot\nabla\phi_{D_{2}}\|_{N-1}\lesssim\lambda_{q+1}^{N}\frac{\rho^{3/2}}{\mu}\lambda_{q}\delta_{q}^{1/2}(\lambda_{q+1}^{-2}+(\lambda_{q+1}\mu)^{-n_{0}-1}).$
This concludes the estimates for $\phi_{D_{2}}$. Regarding the estimates on $f_{D_{2}}$, notice that
(5.110) $\int_{\mathbb{T}^{3}}\left(\mathcal{R}(G_{i}):\nabla u_{l}^{T}\right)(x,s)\,dx=\int_{\mathbb{T}^{3}}\left(\mathcal{R}(\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}):\nabla u_{l}^{T}\right)(x,s)\,dx.$
Therefore, (5.93) implies that
(5.111) $|f_{D_{2}}(t)|+|f^{\prime}_{D_{2}}(t)|\lesssim\frac{\rho}{\mu}\lambda_{q}\delta_{q}^{1/2}(\lambda_{q+1}\mu)^{-n_{0}-1}.$
5.1.3. Proof of Proposition˜5.1: bounds on $\phi_{S}$
Lemma 5.5.
Let $\phi_{S_{1}}$ be defined by (5.26). For $N\leq 2$, it holds for any $t\leq\mathfrak{t}$ that
(5.112) $\|\phi_{S_{1}}(t)\|_{N}$ $\lesssim l^{-N}i_{q}^{1/2-\delta}+\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2}+i_{q}^{1/2-\delta}\begin{cases}1,\quad&N=0\\ \lambda_{q+1}^{N}\delta_{q+1}^{1/2},\quad&N>0\end{cases},$
(5.113) $\|D_{t,q+1}\phi_{S_{1}}(t)\|_{N-1}$ $\lesssim i_{q}^{1/2-\delta}\lambda_{q+1}^{N}\delta_{q+1}+\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2}+\lambda_{q+1}^{N-1}\delta_{q+2}^{3/2}\rho^{1/2}\epsilon^{-1}$
(5.114) $+i_{q+1}^{-1/2-\delta}\begin{cases}1,\quad&N=1\\ \lambda_{q+1}^{N-1}\delta_{q+1}^{1/2}\vee l^{1-N},\quad&N>1\end{cases}$
Proof.
We divide the right-hand side of (5.26) into several terms, each to be controlled separately. First, the trace term can be controlled using the bounds (4.11), (4.12), and (3.60), more specifically we have
(5.115) $\left\|\operatorname{Tr}\left(R_{q+1}-\frac{3}{2}f\,\mathrm{Id}\right)\overline{z}\right\|_{N}$ $\lesssim\lambda_{q+1}^{N}\delta_{q+2}\,i_{q}^{1/2-\delta},$
(5.116) $\left\|D_{t,q+1}\left[\operatorname{Tr}\left(R_{q+1}-\frac{3}{2}f\,\mathrm{Id}\right)\overline{z}\right]\right\|_{N-1}$ $\lesssim\lambda_{q+1}^{N-1}\delta_{q+2}(i_{q+1}^{-1/2-\delta}+i_{q}^{1/2-\delta}\lambda_{q+1}).$
Next, notice that
(5.117) $\frac{|v_{q}|^{2}+|w|^{2}}{2}\overline{z}\,+\,(\overline{z}\otimes v_{q})w=\frac{|v_{q+1}|^{2}}{2}\overline{z},$
for which we have the estimates
(5.118) $\left\|\frac{|v_{q+1}|^{2}}{2}\overline{z}\right\|_{N}$ $\lesssim i_{q}^{1/2-\delta}\begin{cases}1,\quad&N=0\\ \lambda_{q+1}^{N}\delta_{q+1}^{1/2},\quad&N>0\end{cases},$
(5.119) $\left\|D_{t,q+1}\left(\frac{|v_{q+1}|^{2}}{2}\overline{z}\right)\right\|_{N-1}$ $\lesssim\lambda_{q+1}^{N}\delta_{q+1}i_{q}^{1/2-\delta}+i_{q+1}^{-1/2-\delta}\begin{cases}1,\quad&N=0\\ \lambda_{q+1}^{N-1}\delta_{q+1}^{1/2},\quad&N>0\end{cases},$
where the second inequality we have used Lemmas˜3.4, 3.60, 4.11 and 4.40 and
(5.120) $D_{t,q+1}v_{q+1}=\mathord{{\rm div}}\left(R_{q+1}-\frac{3}{2}f\,\mathrm{Id}\right)-\nabla p_{q+1}+u_{q+1}\cdot\nabla z_{q+1}.$
Finally, upon recalling the definition (4.3) of the pressure increment $\overline{p}$, the remaining term $\tilde{\phi}_{S_{1}}:=(\overline{z}\otimes\overline{z}+u_{q}\odot\overline{z})v_{l}+\overline{p}w$ is controlled by means of Proposition˜3.3, Lemma˜3.4, the bounds (2.14), (3.60), (3.51), (4.38), (3.61), and the following inequality:
(5.121) $\|D_{t,q+1}v_{l}\|_{N-1}\leq\|D_{t,l}v_{l}\|_{N-1}+\|(u_{q+1}-u_{l})\cdot\nabla v_{l}\|_{N-1}\lesssim\lambda_{q+1}^{N-1}\lambda_{q}\,\delta_{q}.$
Namely, we have
(5.122) $\|\tilde{\phi}_{S_{1}}\|_{N}$ $\lesssim l^{-N}i_{q}^{1/2-\delta}+\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2},$
(5.123) $\|D_{t,q+1}\tilde{\phi}_{S_{1}}\|_{N-1}$ $\lesssim i_{q+1}^{-1/2-\delta}l^{1-N}+i_{q}^{1/2-\delta}\lambda_{q+1}^{N-1}\lambda_{q}\delta_{q}$
(5.124) $\quad+\lambda_{q+1}^{N-2\gamma}\delta_{q+2}^{3/2}\delta_{q+1}^{1/2}+\lambda_{q+1}^{N-1}\delta_{q+2}^{3/2}\rho^{1/2}\epsilon^{-1}.$
Lemma 5.6.
Let $\phi_{S_{2}}$, $f_{S_{2}}$ be defined by (5.27) and (5.20). For $N\leq 2$, it holds for any $t\leq\mathfrak{t}$ that
(5.125) $\|\phi_{S_{2}}(t)\|_{N}$ $\lesssim\lambda_{q+1}^{N}\delta_{q+2}i_{q}^{1/2-\delta}+i_{q}^{1/2-\delta}\begin{cases}1,\quad&N=0\\ \lambda_{q}^{N}\delta_{q}^{1/2},\quad&N>0\end{cases},$
(5.126) $\|D_{t,q+1}\phi_{S_{2}}(t)\|_{N-1}$ $\lesssim\lambda_{q+1}^{N-1}\lambda_{q}\delta_{q}^{1/2}i_{q}^{1/2-\delta}+\lambda_{q+1}^{N-1}\delta_{q+2}i_{q+1}^{-1/2-\delta}+l\lambda_{q+1}^{N}\delta_{q+2}+i_{q+1}^{-1/2-\delta}\begin{cases}1,\quad&N=1\\ \lambda_{q}^{N-1}\delta_{q}^{1/2},\quad&N>1\end{cases},$
(5.127) $|f_{S_{2}}(t)|+|f^{\prime}_{S_{2}}(t)|$ $\lesssim i_{q}^{1/2-\delta}.$
Proof.
Since we will only use that $\|\overline{\mathcal{R}}\|_{L^{\infty}\to L^{\infty}}\lesssim 1$, the estimates on $f_{S_{2}}$ follow from the estimate on $\|\phi_{S_{2}}\|_{0}$.
Let us start with an estimate on $\bar{\phi}_{S_{2}}:=\overline{\mathcal{R}}[R_{A}:\nabla(z_{q}-z_{l})^{T}]$. In this case, (4.11) implies that
(5.128) $\|\bar{\phi}_{S_{2}}\|_{N}\lesssim l\lambda_{q+1}^{N-\gamma}\delta_{q+2}.$
As for the advective derivative, we have that
(5.129) $\|D_{t,q+1}\bar{\phi}_{S_{2}}\|_{N-1}$ $\leq\left\|\overline{\mathcal{R}}[D_{t,q+1}\left(R_{A}:\nabla(z_{q}-z_{l})^{T}\right)]\right\|_{N-1}+\left\|[u_{q+1}\cdot\nabla,\overline{\mathcal{R}}](R_{A}:\nabla(z_{q}-z_{l})^{T})\right\|_{N-1}$
$\leq\left\|D_{t,q+1}\left(R_{A}:\nabla(z_{q}-z_{l})^{T}\right)\right\|_{N-1}+\sum_{j=0}^{N-1}\|u_{q+1}\|_{j}\|R_{A}:\nabla(z_{q}-z_{l})^{T}\|_{N-j}$
$\leq l\lambda_{q+1}^{N-1}\delta_{q+2}\left(i_{q}^{-1/2-\delta}+\lambda_{q+1}\right),$
where the last inequality is due to Lemma˜3.4 and the estimates (4.11) and (4.12).
The next term we study is $\tilde{\phi}_{S_{2}}:=\overline{\mathcal{R}}[(u_{q}\,\mathring{\odot}\,\overline{z}+\overline{z}\otimes\overline{z}):\nabla z_{q}^{T}]$, which is controlled similarly to bounds (5.122) and (5.123) previously obtained for $\tilde{\phi}_{S_{1}}$, with the difference that (5.121) is now replaced by the bound $\|D_{t,q+1}\nabla z_{q}\|_{N}\lesssim i_{q}^{-1/2-\delta}+\begin{cases}1,\quad&N=0\\ \lambda_{q+1}^{N}\delta_{q+1}^{1/2},\quad&N>0\end{cases}$, consequence of Lemma˜3.4, giving
(5.130) $\|\tilde{\phi}_{S_{2}}\|_{N}\lesssim i_{q}^{1/2-\delta}\begin{cases}1,\quad&N=0\\ \lambda_{q}^{N}\delta_{q}^{1/2},\quad&N>0\end{cases},$
(5.131) $\|D_{t,q+1}\tilde{\phi}_{S_{2}}\|_{N-1}\lesssim\lambda_{q+1}^{N-1}\lambda_{q}\delta_{q}^{1/2}\,i_{q}^{1/2-\delta}+i_{q+1}^{-1/2-\delta}\begin{cases}1,\quad&N=1\\ \lambda_{q}^{N-1}\delta_{q}^{1/2},\quad&N>1\end{cases}.$
Furthermore, similar arguments leading to (5.115) and (5.116) also yield
(5.132) $\left\|\overline{\mathcal{R}}\left[\left(R_{q+1}-\frac{3}{2}f\,\mathrm{Id}\right):\nabla\overline{z}^{T}\right]\right\|_{N}$ $\lesssim\lambda_{q+1}^{N}\delta_{q+2}\,i_{q}^{1/2-\delta},$
(5.133) $\left\|D_{t,q+1}\overline{\mathcal{R}}\left[\left(R_{q+1}-\frac{3}{2}f\,\mathrm{Id}\right):\nabla\overline{z}^{T}\right]\right\|_{N-1}$ $\lesssim\lambda_{q+1}^{N-1}\delta_{q+2}(i_{q+1}^{-1/2-\delta}+i_{q}^{1/2-\delta}\lambda_{q+1}).$
At last, Lemma˜3.4, (4.38), and the following inequalities
(5.134) $\|D_{t,q+1}v_{q}\|_{N}$ $\lesssim\lambda_{q+1}^{N}\lambda_{q}\delta_{q},$
(5.135) $\|D_{t,q+1}\nabla\overline{z}\|_{N}$ $\lesssim i_{q+1}^{-1/2-\delta}+i_{q}^{1/2-\delta}\begin{cases}1,\quad&N=0\\ \lambda_{q+1}^{N}\delta_{q+1}^{1/2},\quad&N>0\end{cases},$
(5.136) $\|D_{t,q+1}\nabla z_{q+1}\|_{N}$ $\lesssim i_{q+1}^{-1/2-\delta}+\begin{cases}1,\quad&N=0\\ \lambda_{q+1}^{N}\delta_{q+1}^{1/2},\quad&N>0\end{cases},$
imply that
(5.137) $\left\|\overline{\mathcal{R}}\Big[v_{q}\cdot\mathord{{\rm div}}(u_{q}\otimes\overline{z}+\overline{z}\otimes z_{q+1})\Big]\right\|_{N}$ $\lesssim i_{q}^{1/2-\delta}\begin{cases}1,\quad&N=0\\ \lambda_{q}^{N}\delta_{q}^{1/2},\quad&N>0\end{cases},$
(5.138) $\left\|D_{t,q+1}\overline{\mathcal{R}}\Big[v_{q}\cdot\mathord{{\rm div}}(u_{q}\otimes\overline{z}+\overline{z}\otimes z_{q+1})\Big]\right\|_{N-1}$ $\lesssim i_{q}^{1/2-\delta}\lambda_{q+1}^{N-1}\lambda_{q}\,\delta_{q}^{1/2}+i_{q+1}^{-1/2-\delta}\begin{cases}1,\quad&N=1\\ \lambda_{q}^{N-1}\delta_{q}^{1/2},\quad&N>1\end{cases}.$
5.1.4. Proof of Proposition˜5.1: Conclusion
Bounds (5.34) and (5.35) follow from Lemma˜5.2, Lemma˜5.3, Lemma˜5.4, Lemma˜5.5, Lemma˜5.6 above and our choice of parameters, up to taking the constant $a_{0}>1$ large enough so to absorb all the implicit constants coming from the aforementioned lemmas. As for the estimates on $f$, we have $f:=f_{A}+f_{D_{2}}+f_{S_{2}}$ and thus by Lemma˜5.2, Lemma˜5.4, and Lemma˜5.6:
(5.139) $|\partial_{t}f(t)|\leq|\partial_{t}f_{A}(t)+\partial_{t}f_{D_{2}}(t)+\partial_{t}f_{S_{2}}(t)|\lesssim\lambda_{q+1}^{-3\gamma/2}\delta_{q+2}\delta_{q+1}^{1/2}.$
Finally, the claim on $f$ follows from $|f(t)|\leq T\sup_{[0,\mathfrak{t}]}|\partial_{t}f(t)|$ and from picking $a_{0}>1$ large enough.
Appendix A Appendix
A.1. Estimates for transport equations
We report some well-known estimates regarding smooth solutions $f:\mathbb{R}^{3}\to\mathbb{R}$ of the transport equation
(A.1) $\begin{cases}&\partial_{t}f+v\cdot\nabla f=g,\\ &f(\cdot,t_{0})=f_{0},\end{cases}$
where $v:\mathbb{R}^{3}\to\mathbb{R}^{3}$ is a smooth vector field and $g:\mathbb{R}^{3}\to\mathbb{R}$ is a smooth scalar field. The following proposition is a slight modification of [BuDLSzVi19, Proposition B.1].
Proposition A.1.
In the same setting as above, for any $\alpha\in[0,1]$ it holds
(A.2) $\|f_{t}\|_{\alpha}\lesssim e^{\alpha(t-t_{0})\|v_{[t_{0},t]}\|_{1}}\left(\|f_{0}\|_{\alpha}+\int_{t_{0}}^{t}\|g_{s}\|_{\alpha}ds\right),\qquad\|v_{[t_{0},t]}\|_{1}:=\sup_{s\in[t_{0},t]}\|v_{s}\|_{C^{1}_{x}}.$
In addition, if $|t-t_{0}|\|v_{[t_{0},t]}\|_{1}\leq 1$, then for any $N\geq 1$ and $\alpha\in[0,1)$
(A.3) $f_{t}]_{N+\alpha}\lesssim[f_{0}]_{N+\alpha}+|t-t_{0}|[v_{[t_{0},t]}]_{N+\alpha}[f_{0}]_{1}+\int_{t_{0}}^{t}\left([g_{s}]_{N+\alpha}+(t-s)[v_{[t_{0},t]}]_{N+\alpha}[g_{s}]_{1}\right)ds,$
where $v_{[t_{0},t]}]_{N+\alpha}:=\sup_{s\in[t_{0},t]}[v_{s}]_{N+\alpha}$.
Moreover, the spatial inverse $\phi$ of the Lagrangian flux:
(A.4) $\phi_{t}(x)=(X_{t})^{-1}(x),\quad\dot{X}_{t}=v(t,X_{t}),\quad X(x,t_{0})=x,\quad\forall x\in\mathbb{R}^{3},$
satisfies the bounds
(A.5) $\|\nabla\phi(t)-\mathrm{Id}\|_{0}$ $\lesssim|t-t_{0}|\|v_{[t_{0},t]}\|_{1},$
(A.6) $[\phi(t)]_{N}$ $\lesssim|t-t_{0}|[v_{[t_{0},t]}]_{N},\quad\forall N\geq 2.$
A.2. Antidivergence estimates
In this subsection, we report a technical generalization of [DLK23, Corollary 8.2]. Let us start by recalling the following lemma from [DLK23].
Lemma A.2.
([DLK23, Lemma 8.1 ]). Let $\mathfrak{m}\in\mathcal{S}(\mathbb{R}^{3})$ and define the Fourier multiplier operator $T_{\mathfrak{m}}$, acting an smooth functions $f\in C^{\infty}(\mathbb{T}^{3},\mathbb{C})$ (smoothness meant in real sense) as follows:
(A.7) $\mathcal{F}[T_{\mathfrak{m}}(f)](k)=\mathcal{F}[f](k)\mathfrak{m}(k),\quad k\in\mathbb{Z}^{3}.$
Then, for any $n_{0}\in\mathbb{N}$, $\lambda>0$ and any scalar functions $a,\,\xi\in C^{\infty}(\mathbb{T}^{3})$, the function $T_{\mathfrak{m}}(ae^{i\lambda\xi})$ can be decomposed as
(A.8) $T_{\mathfrak{m}}(ae^{i\lambda\xi})=\left[a\,\mathfrak{m}(\lambda\nabla\xi)+\sum_{k=1}^{2n_{0}}C_{k}^{\lambda}(\xi,a):(\nabla^{k}\mathfrak{m})(\lambda\nabla\xi)+\mathcal{\epsilon}_{n_{0}}(\xi,a)\right]e^{i\lambda\xi},$
for some tensor-valued coefficents $C_{k}^{\lambda}(\xi,a)$ and smooth function $\mathcal{\epsilon}_{n_{0}}$ given by
(A.9) $\mathcal{\epsilon}_{n_{0}}(\xi,a)(x)$ $:=\sum_{n_{1}+n_{2}=n_{0}+1}\frac{(-1)^{n_{1}}c_{n_{1},n_{2}}}{n_{0}!}$
(A.10) $\quad\cdot\int_{0}^{1}\int_{\mathbb{R}^{3}}\overset{\vee}{\mathfrak{m}}(y)e^{-i\lambda\nabla\xi(x)\cdot y}((y\cdot\nabla)^{n_{1}}a)(x-ry)e^{i\lambda Z[\xi]_{x,y}(r)}\beta_{n_{2}}[\xi](r)(1-r)^{n_{0}}dy\,dr,$
where $c_{n_{1},n_{2}}$ are coefficients depending only on $n_{1},n_{2}$, and $\beta_{n}[\xi]$ is defined by the following expressions:
(A.11) $\beta_{n}[\xi](r):=B_{n}(i\lambda Z^{\prime}(r),\cdots,i\lambda Z^{(n)}(r)),$
(A.12) $Z(r):=Z[\xi]_{x,y}(r):=r\int_{0}^{1}(1-s)(y\cdot\nabla)^{2}\xi(x-rsy)\,ds,\quad x,y\in\mathbb{R}^{3},$
(A.13) $B_{n}(x_{1},\dots,x_{n}):=\sum_{k=1}^{n}B_{n,k}(x_{1},x_{2},\dots,x_{n-k+1}),$
(A.14) $B_{n,k}(x_{1},x_{2},\dots,x_{n-k+1}):=\sum\frac{n!}{j_{1}!j_{2}!\cdots j_{n-k+1}!}\left(\frac{x_{1}}{1!}\right)^{j_{1}}\cdots\left(\frac{x_{n-k+1}}{n-k+1!}\right)^{j_{n-k+1}},$
and the last summation is over $\{j_{k}\}\subset\mathbb{N}$ such that
(A.15) $j_{1}+\cdots+j_{n-k+1}=k,\quad j_{1}+2j_{2}+\cdots+(n-k+1)j_{n-k+1}=n.$
Proposition A.3.
Let $v\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$ and $l,\epsilon,\lambda$ be positive constants. Let $v_{l}:=\mathbf{P}_{\leq l^{-1}}v$ and consider, for any $m\in\mathbb{Z}$, the solution $\xi_{m}$ to the vector-valued transport equation
(A.16) $\begin{cases}&\partial_{t}\xi_{m}+v_{l}\cdot\nabla\xi_{m}=0,\\ &\xi_{m}(x,\epsilon(m-1/8))=x,\quad\forall x\in\mathbb{R}^{3}.\end{cases}$
Let $F:=\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\sum_{m\in\mathbb{Z}}a_{m,k}e^{i\lambda\xi_{m}\cdot k}$, where the functions $a_{m,k}$ satisfy the following conditions:
  1. (1)
    $\operatorname{Supp}_{t}(a_{m,k})\subset(\epsilon(m-1/8),\epsilon(m+9/8))$, and $\epsilon>0$ is such that $2\epsilon\|v\|_{C_{t}C^{1}_{x}}\leq 1/4$;
  2. (2)
[…]
  3. (3)
    the sequence $(s_{n})_{n\in\mathbb{N}}$ is non-decreasing and the following inequalities hold for every $n\leq m$:
    (A.18) $l^{-m}s_{n}\leq s_{m+n},\quad s_{n}\lambda^{m}\leq s_{0}\lambda^{n+m},\quad s_{m}\lambda^{-m}\leq s_{n}\lambda^{-n}.$
Then, for every $N\in\{0,1,2\}$ and every $t\in\mathbb{R}$ we have
(A.19) $\|\mathcal{R}F(t)\|_{N}\lesssim\left(s_{0}\lambda^{N-1}+\lambda^{-n_{0}-1}s_{n_{0}+1}\lambda^{N}\right)a_{F}.$
Proof.
First of all, let us quickly show (more details can be found at [DLK23]) that
(A.20) $F=\mathbf{P}_{\gtrsim 3\lambda/8}\left(\sum_{m,k}a_{m,k}e^{i\lambda k\cdot\xi_{m}}\right)-\sum_{m,k}\epsilon_{n_{0}}^{\lambda}(k\cdot\xi_{m},a_{m,k})e^{i\lambda k\cdot\xi_{m}},$
where
(A.21) $\mathbf{P}_{\gtrsim 3\lambda/8}$ $=\sum_{2^{j}\geq 3\lambda/8}\mathbf{P}_{2^{j}},$
(A.22) $\epsilon_{n_{0}}^{\lambda}(k\cdot\xi_{m},a_{m,k})$ $=\sum_{2^{j}\geq 3\lambda/8}\epsilon_{n_{0},j}(k\cdot\xi_{m},a_{m,k}),$
with $\epsilon_{n_{0},j}$ being obtained by Lemma˜A.2 applied to the Fourier multiplier operator $\mathbf{P}_{2^{j}}$. Notice that, by construction, $\xi_{m}$ is smooth and space-periodic, since $v_{l}$ is so.
In order to show (A.20), let us consider $(m,k)\in\mathbb{Z}\times(\mathbb{Z}^{3})^{*}$ and $j_{k}\in\mathbb{N}$ such that $2^{j_{k}}\geq\frac{5}{4}\lambda|k|$. Then
(A.23) $\mathbf{P}_{\gtrsim 3\lambda/8}=\sum_{3\lambda/8\leq 2^{i}\leq 2^{j_{k}}}\mathbf{P}_{2^{i}}+\mathbf{P}_{>2^{j_{k}}}=:T_{\mathfrak{m}_{1}}+\sum_{j>j_{k}}T_{\mathfrak{m}_{j}},$
where the functions $\mathfrak{m}_{1}\in\mathcal{S}(\mathbb{R}^{3})$ and $T_{\mathfrak{m}_{j}}=\mathbf{P}_{2^{j}}$ satisfy
(A.24) $\operatorname{Supp}\,\mathfrak{m}_{1}\subset B(0,2^{j_{k}+1}),\,\quad\mathfrak{m}_{1}\equiv 1\,\text{ on }\,\overline{B(0,2^{j_{k}})}\setminus B(0,\frac{3}{4}\lambda|k|),\quad\mathfrak{m}_{j}\equiv 0\text{ on }\overline{B}(0,\frac{5}{4}\lambda|k|).$
Moreover for every $t\in\operatorname{Supp}_{t}a_{m,k}$, Assumption 1 of this proposition and Eq.˜A.5 imply that
(A.25) $\||(\nabla\xi_{m}(t))^{T}k|-|k|\|_{0}\leq 2\epsilon|k|\|v\|_{C_{t}C_{x}^{1}}\leq\frac{|k|}{4}\implies\|\nabla(k\cdot\xi_{m}(t))\|_{0}\in\left[\frac{3}{4}|k|,\frac{5}{4}|k|\right]$
Therefore, Lemma˜A.2 applied to $T_{m_{1}}$, and noticing that $\mathfrak{m}_{1}(\lambda\nabla(k\cdot\xi_{m}))\equiv 1$, $\nabla^{h}\mathfrak{m}_{1}(\lambda\nabla(k\cdot\xi_{m}))\equiv 0$ for every $h\geq 1$, giving:
(A.26) $T_{\mathfrak{m}_{1}}(a_{m,k}e^{i\lambda k\cdot\xi_{m}})$ $-\sum_{3\lambda/8\leq 2^{i}\leq 2^{j_{k}}}\mathcal{\epsilon}_{n_{0},i}(k\cdot\xi_{m},a_{m,k})e^{i\lambda k\cdot\xi_{m}}$
(A.27) $=a_{m,k}\,\underbrace{\mathfrak{m}_{1}(\lambda\nabla(k\cdot\xi_{m}))}_{\equiv 1}e^{i\lambda k\cdot\xi_{m}}+\sum_{h=1}^{2n_{0}}C_{h}^{\lambda}(k\cdot\xi_{m},a_{m,k}):(\underbrace{\nabla^{h}\mathfrak{m}_{1}}_{\equiv 0})(\lambda\nabla(k\cdot\xi_{m}))e^{i\lambda k\cdot\xi_{m}}$
(A.28) $=a_{m,k}e^{i\lambda k\cdot\xi_{m}}.$
At last, Eq.˜A.20 follows from applying Lemma˜A.2 to $T_{\mathfrak{m}_{j}}$ for all $j>j_{k}$. Indeed,
(A.29) $T_{\mathfrak{m}_{j}}(a_{m,k}e^{i\lambda k\cdot\xi_{m}})=\epsilon_{n_{0},j}(k\cdot\xi_{m},a_{m.k})e^{i\lambda k\cdot\xi_{m}}$
thanks to $(\nabla^{h}\mathfrak{m}_{j})(\lambda\nabla(k\cdot\xi_{m}))=0$ for all $h\in\mathbb{N}$.
Now, for fixed $(m,k)$, we want to estimate $\|\epsilon_{n_{0}}^{\lambda}(k\cdot\xi_{m},a_{m,k})\|_{0}$. Based on Eq.˜A.22, we have to consider $\epsilon_{n_{0},j}(k\cdot\xi_{m},a_{m.k})$, which is given by Eq.˜A.9.
Let $Z$ be defined by Eq.˜A.12 and $Z_{0}(r):=r^{-1}Z(r)$, then
(A.30) $Z^{(n)}=nZ_{0}^{(n-1)}+Z_{0}^{(n)},\quad Z^{(n)}(r):=\int_{0}^{1}(1-s)(-s)^{n}(y\cdot\nabla)^{n+2}(k\cdot\xi_{m})(x-rsy)\,ds.$
Therefore, Assumption 1 and Proposition˜A.1 imply that
(A.31) $\|Z^{(n)}\|_{0}\lesssim|y|^{n+1}\|k\cdot\xi_{m}\|_{n+1}+|y|^{n+2}\|k\cdot\xi_{m}\|_{n+2}\lesssim|y|^{n+1}l^{-n}|k|[1+|y|l^{-1}].$
Then, by means of Eqs.˜A.13 and A.14, we have that
(A.32) $\|\beta_{n_{2}}[k\cdot\xi_{m}](r)\|_{0}$ $\lesssim\sum_{m=1}^{n_{2}}\lambda^{m}|y|^{m+n_{2}}|k|^{m}[1+|y|l^{-1}]^{m}l^{-n_{2}}$
(A.33) $\lesssim(|y|l^{-1})^{n_{2}}|k|^{n_{0}+1}\sum_{m=1}^{n_{2}}(\lambda|y|)^{m}\sum_{i=0}^{m}(|y|l^{-1})^{i}.$
Instead, Assumption 2 implies that
[…]
Then, Eqs.˜A.32 and A.34 imply that
[…]
where the last inequality is due to $\|\overset{\vee}{\mathfrak{m}}_{j}(y)|y|^{n}\|_{L^{1}}\lesssim 2^{-jn}$. Therefore, we have that
[…]
where the last inequality makes use of Assumption 3. The assumption on the time supports, i.e. condition 1, implies that
[…]
So, Eq.˜A.20 allows to conclude as follows:
(A.39) $\|\mathcal{R}F\|_{N}$ $\leq\|\mathcal{R}\mathbf{P}_{\gtrsim 3\lambda/8}F\|_{N}+\|\mathcal{R}\sum_{m,k}\epsilon^{\lambda}_{n_{0}}(k\cdot\xi_{m},a_{m,k})\|_{N}$
(A.40) $\lesssim\frac{\|F\|_{N}}{\lambda}+\lambda^{N}\|\sum_{m,k}\epsilon_{n_{0}}^{\lambda}(k\cdot\xi_{m},a_{m,k})\|_{0}$
(A.41) $\lesssim a_{F}[s_{0}\,\lambda^{N-1}+s_{n_{0}+1}\lambda^{-n_{0}-1}],$
where the second inequality is due to the Bernstein inequality, while the third one is due to
[…]
and the last inequality holds because of Assumption 3. ∎
Proposition A.4.
Under the same assumptions of Proposition˜A.3, let $l>\frac{100}{\lambda}$ and $w\in[C^{\infty}(\mathbb{T}^{3})]^{3}$ such that there exists a non-decreasing sequence $(b_{n})_{n\in\mathbb{N}}$ satisfying for every $n,m\in\mathbb{N}$
(A.43) $b_{n}\lambda^{m}\leq b_{0}\lambda^{m+n},\quad\|w\|_{n}\leq b_{n}.$
Let us moreover suppose there exists $c>0$ such that
[…]
where $D_{t,l}:=\partial_{t}+v_{l}\cdot\nabla$ and $j\leq(n_{0}+1)\vee 2$, and that for any $n\in\mathbb{N}$
(A.45) $\|v\|_{n+1}\leq l^{-n}\|v\|_{1}.$
Then for every $N\leq 1$ and $h_{0}\in\mathbb{N}\setminus\{0,1\}$ it holds
(A.46) $\|D_{t}\mathcal{R}F(t)\|_{N}$ $\lesssim\Big[s_{0}(c\lambda)^{-1}+c^{-1}s_{n_{0}+1}\lambda^{-n_{0}-1}+s_{0}\lambda^{-1}\|v\|_{1}$
(A.47) $\quad+(1+\lambda^{-n_{0}}s_{n_{0}+1})(b_{0}+l\|v\|_{1}+\|v\|_{0})+s_{0}(l\lambda)^{-h_{0}}\lambda^{2}\|v\|_{1}\Big]\lambda^{N}a_{F},$
where $D_{t}:=\partial_{t}+(w+v)\cdot\nabla$.
Proof.
By standard computations, we have that
(A.48) $D_{t}\mathcal{R}F=\mathcal{R}(D_{t,l}F)+[v_{l}\cdot\nabla,\mathcal{R}]F+[(w+(v-v_{l}))\cdot\nabla]\mathcal{R}F.$
Since $D_{t,l}F=\sum_{k,m}D_{t,l}a_{m,k}e^{i\lambda k\cdot\xi_{m}}$, then the previous Proposition˜A.3 implies that
(A.49) $\|\mathcal{R}(D_{t,l}F)\|_{N}\lesssim a_{F}\lambda^{N}[s_{0}(c\lambda)^{-1}+s_{n_{0}+1}\lambda^{-n_{0}-1}c^{-1}].$
Instead,
(A.50) $\|[(w+(v-v_{l}))\cdot\nabla]\mathcal{R}F\|_{N}$ $\lesssim\sum_{N_{1}+N_{2}\leq N}\|w+(v-v_{l})\|_{N_{1}}\|RF\|_{N_{2}+1}$
(A.51) $\lesssim\sum_{N_{1}+N_{2}\leq N}(b_{N_{1}}+\|v\|_{1}l^{1-N_{1}})\lambda^{N_{2}}(1+s_{n_{0}+1}\lambda^{-n_{0}})$
(A.52) $\lesssim(b_{0}+l\|v\|_{1})(1+s_{n_{0}+1}\lambda^{-n_{0}})\lambda^{N},$
where the second inequality is due to the previous result, Bernstein inequality, (A.45) and the assumptions on $w$, while the third is due to $b_{n}\lambda^{m}\leq b_{0}\lambda^{m+n}$.
At last, we need to estimate
(A.53) $v_{l}\cdot\nabla,\mathcal{R}]F=[v_{l}\cdot\nabla,\mathcal{R}]\mathbf{P}_{\gtrsim\lambda}F-[v_{l}\cdot\nabla,\mathcal{R}]\sum_{m,k}\epsilon^{\lambda}_{n_{0}}(k\cdot\xi_{m},a_{m,k})e^{i\lambda k\cdot\xi_{m}}=:I_{1}+I_{2}.$
Since $I_{2}$ has localized active frequencies, the Bernstein inequality implies that
(A.54) $\|I_{2}\|_{N}\lesssim\lambda^{N}\|I_{2}\|_{0}$ $\lesssim\lambda^{N}\|v_{l}\|_{0}\|\nabla\mathbf{P}_{\lesssim\lambda}F\|_{0}$
$\lesssim\lambda^{N+1}\|v\|_{0}\|\sum_{m,k}\epsilon^{\lambda}_{n_{0}}(k\cdot\xi_{m},a_{m,k})\|_{0}$
$\lesssim\lambda^{N}\|v\|_{0}s_{n_{0}+1}\lambda^{-n_{0}}a_{F},$
where the last inequality is due to Eq.˜A.38. Instead, by considering $\mathcal{R}$ as a Fourier multiplier with coefficients $\mathcal{F}(\mathcal{R})(k)$, see [RT25] for an explicit formula, we have that
(A.55) $-I_{1}$ $=\sum_{2^{j}\gtrsim\lambda}\sum_{k,n}i\eta e^{ik\cdot x}\mathcal{F}(v_{l})(k-\eta)\mathcal{F}[F_{j}](\eta)[\mathcal{F}(\mathcal{R})(k)-\mathcal{F}(\mathcal{R})(\eta)]$
$=\sum_{2^{j}\gtrsim\lambda}\sum_{k,n}i\eta\mathcal{F}(v_{l})(k-\eta)\mathcal{F}(F_{j})(\eta)e^{ik\cdot x}\Big[\sum_{h=1}^{h_{0}}\frac{1}{h!}[(k-\eta)\cdot\nabla_{\eta}]^{h}\mathcal{F}(\mathcal{R})(\eta)$
$+\frac{1}{h_{0}!}\int_{0}^{1}[(k-\eta)\cdot\nabla_{\eta}]^{h_{0}+1}\mathcal{F}(\mathcal{R})(\eta+\sigma(k-\eta))(1-\sigma)^{h_{0}}d\sigma\Big]$
$=:J_{1}+J_{2},$
where $F_{j}=\mathbf{P}_{2^{j}}F.$ Now,
(A.56) $\|J_{1}\|_{N}$ $\lesssim\sum_{h=1}^{h_{0}}\|\nabla^{h}v_{l}\nabla\mathcal{R}_{h}\mathbf{P}_{\gtrsim\lambda}F\|_{N}\lesssim\sum_{h=1}^{h_{0}}\sum_{N_{1}+N_{2}\leq N}l^{1-h-N_{1}}\|v\|_{1}\|F\|_{N_{2}+1}\lambda^{-h-1}$
$\lesssim s_{0}\|v\|_{1}\lambda^{-2}\sum_{N_{1}+N_{2}\leq N}l^{-N_{1}}\lambda^{N_{2}+1}a_{F}\lesssim a_{F}s_{0}\|v\|_{1}\lambda^{N-1},$
where $\mathcal{R}_{h}g$ has Fourier coefficient at frequency $k\in\mathbb{Z}^{3}\setminus\{0\}$ given by $\left[\nabla^{h}\mathcal{F}(\mathcal{R})\right](k)\mathcal{F}(g)(k)$, and the second-to-last inequality is due to Eq.˜A.42.
At last,
(A.57) $\|J_{2}\|_{N}$ $\lesssim\sum_{2^{j}\gtrsim\lambda}\sum_{k,\eta}|k|^{N}|\eta||\mathcal{F}(v_{l})(k-\eta)||\mathcal{F}(F_{j})(\eta)||k-\eta|^{h_{0}+1}|\eta-\sigma(\eta-k)|^{-h_{0}-2}$
$\lesssim\sum_{2^{j}\gtrsim\lambda}\sum_{k,\eta}|\eta|^{N-h_{0}-1}|\mathcal{F}(F_{j})(\eta)||k-\eta|^{h_{0}}|\mathcal{F}(\nabla v_{l})(k-\eta)|$
$\lesssim\sum_{2^{j}\gtrsim\lambda}\frac{l^{-h_{0}}}{2^{j(h_{0}+1)}}\sum_{|k|\leq 2^{j}}\|\nabla^{N}F_{j}\|_{L^{2}}\|\nabla v\|_{L^{2}}\lesssim(l\lambda)^{-h_{0}}\lambda^{2}\|v\|_{1}s_{0}\,a_{F}\lambda^{N}.$
where the second inequality is due to $l^{-1}<\frac{\lambda}{100}$ and $|k-\eta|\leq 2l^{-1}$, which imply that $|\eta-\sigma(\eta-k)|\geq\frac{|\eta|}{2}$ and $|k|\lesssim|\eta|$. All in all, Eqs.˜A.49, A.50, A.54, A.56 and A.57 prove the claim. ∎
Remark A.5.
Notice that
(A.58) $\sum_{h=1}^{h_{0}}\|\nabla^{h}v_{l}\nabla\mathcal{R}_{h}\mathbf{P}_{\gtrsim\lambda}F\|_{N}\lesssim\sum_{h=1}^{h_{0}}\sum_{N_{1}+N_{2}\leq N}l^{1-h}\|\nabla v_{l}\|_{N_{1}}\lambda^{-h}\|F\|_{N_{2}}.$
This inequality and Eq.˜A.57 imply that
(A.59) $\|[v_{l}\cdot\nabla,\mathcal{R}]\mathbf{P}_{\gtrsim\lambda}F\|_{N}\lesssim(l+(l\lambda)^{-h_{0}}\lambda^{2})\sum_{N_{1}+N_{2}\leq N}\|\nabla v_{l}\|_{N_{1}}\|F\|_{N_{2}}.$
Proposition A.6.
Under the same assumption of Propositions˜A.3 and A.4, it holds that
(A.60) $\|D_{t,l}\sum_{m,k}\epsilon^{\lambda}_{n_{0}}(k\cdot\xi_{m},a_{m,k})e^{i\lambda k\cdot\xi_{m}}\|_{0}\lesssim a_{F}(c^{-1}\vee\|v_{l}\|_{1})s_{n_{0}+1}\lambda^{-n_{0}-1},$
Proof.
For fixed $(m,k)$, we want to estimate $\|D_{t,l}\epsilon_{n_{0}}^{\lambda}\|_{0}$, given by Eq.˜A.22. we have to consider $\epsilon_{n_{0},j}(k\cdot\xi_{m},a_{m.k})$, which is given by Eq.˜A.9. Based on Eq.˜A.30 and $Z_{0}(r):=r^{-1}Z(r)$, we have that
(A.61) $\|D_{t,l}Z^{(n)}_{0}\|_{0}$ $\lesssim\|\int_{0}^{1}(1-s)(-s)^{n}[v_{l}(x)-v_{l}(x-yrs)]\cdot\nabla_{x}[(y\cdot\nabla)^{n+2}(k\cdot\xi_{m})(x-rsy)]ds\|_{0}$
$+\|\int_{0}^{1}(1-s)(-s)^{n}[D_{t,l}(y\cdot\nabla)^{n+2}]k\cdot\xi_{m}(x-rsy)ds\|_{0}$
$\lesssim\|v_{l}\|_{1}|y|^{n+2}l^{-n-1}|k|[1+l^{-1}|y|],$
where the last inequality is due to the following one, which is proved in [DLK23],
(A.62) $\|D_{t,l}(y\cdot\nabla)^{n}g\|_{0}\lesssim|y|^{n}\left[\|D_{t,l}g\|_{n}+\sum_{\begin{subarray}{c}m_{1}+m_{2}=n\vskip 0.1pt\\ m_{1}\geq 1\end{subarray}}\|v_{l}\|_{m_{1}}\|g\|_{m_{2}+1}\right].$
Then, similarly to Eqs.˜A.31 and A.32, we have that
(A.63) $\|D_{t,l}Z^{(n)}\|_{0}$ $\lesssim|y|^{n+1}l^{-n}|k|\|v_{l}\|_{1}[1+|y|l^{-1}+|y|^{2}l^{-2}],$
(A.64) $\|D_{t,l}\beta_{n_{2}}[k\cdot\xi_{m}]\|_{0}$ $\lesssim\|v_{l}\|_{1}[1+|y|^{2}l^{-2}](|y|l^{-1})^{n_{2}}|k|^{n_{0}+1}\sum_{m=1}^{n_{2}}(\lambda|y|)^{m}\sum_{i=0}^{m}(|y|l^{-1})^{i}.$
Now, Eqs.˜3.73 and A.1 imply that
(A.65) $\|D_{t,l}(e^{-i\lambda y\cdot\nabla(k\cdot\xi_{m})})\|_{0}$ $\lesssim\lambda|y|\|v_{l}\|_{1}|k|,$
(A.66) $\|D_{t,l}e^{iZ(k\cdot\xi_{m})}\|_{0}$ $\lesssim\lambda\|\int_{0}^{1}(1-s)(y\otimes y):[(D_{t,l}\nabla^{2}(k\cdot\xi_{m})(x-ry)$
(A.67) $\quad+\left((v_{l}(x)-v_{l}(x-ry))\cdot\nabla\right)\nabla^{2}(k\cdot\xi_{m})(x-ry)ds]e^{i\lambda Z(r)}\|_{0}$
(A.68) $\lesssim\lambda|y|^{2}|k|l^{-1}\|v_{l}\|_{1}[1+|y|l^{-1}].$
Instead, Assumption 2 and Eq.˜A.62 imply that
[…]
By following the computations used to get Eq.˜A.37, Eqs.˜A.64, A.66, A.65 and A.69 imply that
[…]
Then, the claim follows by just summing up in $m$ and $k$. ∎